### v1

# Sommerfeld Enhancement for Vector Dark Matter in the Fundamental Representation of $SU(2)_{L}$

Abstract
In this work, we study an extension of the Standard Model which include a new massive vector field in the fundamental representation of $SU(2)_{L}$. The neutral component of the new vector field is a natural dark matter candidate. We compute the Annihilation Cross Section with Sommerfeld Enhancement and the Photon Flux arising from Dark Matter annihilation events. We found that the model exhibits a resonant mass part around the range of $5-10$ TeV. This resonance is influenced by the choice of portal Higgs coupling parameter. We show that these prediction can be partially tested by CTA in the near future
1 Introduction
Currently, one the most urgent problems in Particle Physics revolves around the elucidation of Dark Matter’s (DM) [2] fundamental nature. It is well known that,to account for the overarching phenomenology of DM, it suffices to posit the existence of a new particle falling within a mass range spanning from a few GeV to a few TeV, characterized by an interaction strength with the standard sector on the order of the weak interaction. This phenomenon is commonly referred to as the ”WIMP miracle,” with WIMP representing ”Weakly Interacting Massive Particle [12].
An intriguing realization of the WIMP idea is to postulate the DM as the neutral component of an electroweak multiplet. This paradigm, usually called “Minimal Dark Matter”, was proposed long ago [8] and have been studied in datail for scalar field [7] and fermionic fields [10].
In the case of vector DM (VDM), an effective model including a VDM candidate originated from a vector field in the adjoint representation of $SU(2)_{L}$ was first proposed in [4] while an ultraviolet complete model was developed in [1]. On the other had, in the fundamental representation, models including a vector DM candidate coming from an electroweak doublet were studied in  [15, 18]. These vector models are consistent with experimental data, if the new fields has mass in the range of some few TeV. In this mass range, direct detection [19] becomes difficult and the discovery at colliders seems to be challenging even for future colliders [4].
In this circumstances, indirect search can be a privileged alternative to discover or rule out these models. However, it is well known that to obtain an accurate value of the DM annhihilation cross section in the non-relativistic regime, it is necessary to take into account non-perturbative effects like the so called Sommerfeld Enhancement (SE).
The SE effect was computed for the cases of scalar DM field [3, 7, 9, 11, 13] showing that effect has a resonant part, fermion DM field [5, 10] in its doublet and triplet representation of $SU(2)_{L}$ and vector DM in the adjoint representation in  [1]. In this work, we focus on computing the SE for the annihilation of VDM in the fundamental representation of $SU(2)_{L}$ and the photon flux (relevant for Indirect Detection  [6, 14, 19]) produced by pair annihilation of the VDM.
Our work is organized as follows. In section 2, we describe the main aspects of the model. Section 3 is devoted to the computation of the annihilation matrices. In section 4 we apply our results to compute the photon flux from DM annihilation at the center of the Milky Way and describe our results. Finally, in section 5 we state our conclusions.
2 The Model
We start by recalling the main characteristics of the model initially developed in reference  [15]. In this context, the Standard Model (SM) is extended by the introduction of a vector field $V_{\mu}$ which transforms as (1,2,1/2) under the standard group $SU(3)_{c}\times SU(2)_{L}\times U(1)_{Y}$ This new vector field can be explicitly written as:
$V_{\mu}=\begin{pmatrix}V_{\mu}^{+}\\ V_{\mu}^{0}\end{pmatrix}=\begin{pmatrix}V_{\mu}^{+}\\ \dfrac{V_{\mu}^{1}+i~{}V_{\mu}^{2}}{\sqrt{2}}\end{pmatrix}$ (5)
In principle, the two neutral states $V_{\mu}^{1}$ and $V_{\mu}^{2}$ (the vector and axial-vector fields) can have different masses. However, the requirement of unitarity preservation at low energies constrains the parameter space of the model in such a way that $V_{\mu}^{1}$ and $V_{\mu}^{2}$ become degenerated or nearly degenerated. In this work, we assume that $V_{\mu}^{1}$ and $V_{\mu}^{2}$ are exactly degenerated. In this case they combine into a complex neutral field $V_{\mu}^{0}$.
The Lagrangian of the model can be written as:
$\mathcal{L}=\mathcal{L}_{SM}$ $-\dfrac{1}{2}\left(D_{\mu}V_{\nu}-D_{\nu}V_{\mu}\right)^{\dagger}\left(D^{\mu}V^{\nu}-D^{\nu}V^{\mu}\right)+M_{V}^{2}V_{\nu}^{\dagger}V^{\nu}$
$-\lambda_{2}(\phi^{\dagger}\phi)(V_{\mu}^{\dagger}V^{\mu})-\lambda_{3}(\phi^{\dagger}V_{\mu})(V^{\mu\dagger}\phi)$
$+\lambda_{4}\left[(\phi^{\dagger}V_{\mu})(\phi^{\dagger}V^{\mu})+(V^{\mu\dagger}\phi)(V_{\mu}^{\dagger}\phi)\right]$
$-\alpha_{2}(V_{\mu}^{\dagger}V^{\mu})(V_{\nu}^{\dagger}V^{\nu})-\alpha_{3}(V_{\mu}^{\dagger}V^{\nu})(V_{\nu}^{\dagger}V^{\mu})$
$+i\dfrac{g^{\prime}}{2}\kappa_{1}V_{\mu}^{\dagger}B^{\mu\nu}V_{\nu}+ig\kappa_{2}V_{\mu}^{\dagger}W^{\mu\nu}V_{\nu}$ (6)
where $B^{\mu\nu}$ represents the Abelian strength field of $U(1)_{Y}$ and $W^{\mu\nu}=W^{\mu\nu~{}a}\dfrac{\tau^{a}}{2}$ represents the non-Abelian strength field of $SU(2)_{L}$. Notice that the las two terms of Lagrangian (2) describe non-minimal interactions between the new vector field and the gauge bosons. In order to avoid a direct photon interaction with $V_{\mu}^{0}$, the otherwise arbitrary constants $\kappa_{1}$ and $\kappa_{2}$ have to be equal. In this work, however, for the sake of simplicity and minimality we assume that $\kappa_{1}=\kappa_{2}=0$. Besides, the coupling constant that regulates the interactions between the Higgs field and the new vector field can be treated a free parameters. For simplicity, we will take the assumption $\lambda_{2}=\lambda_{3}=\lambda_{4}=\lambda$). In this way, $\lambda$ will be the only constant regulating the Higgs portal.
Due to a fortunate conspiracy between the Lorentz and gauge structures of the model, the new vector field can only appear in pairs. That means that the Lagrangian is invariant by a $Z_{2}$ transformation under which the new vector is odd and all the standard fields are even. Consequently, this accidental $Z_{2}$ symmetry makes the lighter component of $V_{\mu}$ stable. This stable component happens to be the neutral one ($V_{\mu}^{0}$)
3 Sommerfeld Enhancement Computation
3.1 Basic Formalism
The so called “Sommerfeld enhancement” is a non-perturbative effect due to the resumation of initial states interactions. This effect may produce sensible consequences for the annihilation cross section in the non-relativistic regime  [10, 17]. In the few next paragraphs we will review the formalism used to compute the Sommerfeld enhancement effect and the we will apply it to our model.
We start by considering the DM annihilation processes, which are described by a two-body state, in the non-relativistic regime. The non-relativistic Schwinger-Dyson equation for the Green function $\bra{0}T\phi_{b}(R^{\prime},\textbf{r}^{\prime})\phi_{a}^{\dagger}(R,\textbf{r})\ket{0}$ can be written as:
$\left[\left(i\partial_{R^{0}}+\frac{\nabla_{R}^{2}}{4m}+\frac{\nabla_{r}^{2}}{m}\right)-V(r)+i\frac{N}{2}\Gamma\delta^{3}(r)\right]$ $\bra{0}T\phi_{b}(R^{\prime},\textbf{r}^{\prime})\phi_{a}^{\dagger}(R,\textbf{r})\ket{0}$ (7)
$=i\delta^{4}(R-R^{\prime})\delta^{3}(\textbf{r}-\textbf{r}^{\prime})\delta_{ab}$
Here, $m$ represents the mass of Dark Matter, $\phi_{i}$ represents the DM field, $N$ is the number spin states for the two bodies system (for spin 1, we have $N=9$ ) and $\Gamma$ is the so called annihilation matrix. The coordinate $=(R^{0},\textbf{R})$ R ) denotes the center-of-mass coordinate, and r represents the relative coordinate of the Dark Matter multiplets, defined as:
[…]
The imaginary part of the diagonal component of this Green function $\bra{0}T\phi_{b}\phi_{a}^{\dagger}\ket{0}$, is associated with the total annihilation cross section. We proceed to decompose the Green functions into individual partial wave modes
$\bra{0}T\phi_{b}(R^{\prime},\textbf{r}^{\prime})\phi_{a}^{\dagger}(R,\textbf{r})\ket{0}=\int\dfrac{d^{4}P}{(2\pi)^{4}}e^{-iP\cdot(R-R^{\prime})}\sum_{l}\dfrac{2l+1}{4\pi}P_{l}(\cos\gamma)(-i)G_{ba}^{(E,l)}(r,r^{\prime})$ (9)
where $\gamma$ denotes the relative angle between r and $\textbf{r}^{\prime}$. The non-relativistic energy for a dark matter pair ($E$) is
$E=P^{0}-\dfrac{\textbf{P}}{4m}=\dfrac{mv^{2}}{4}$ (10)
where, $v$ represents the relative velocity between a pair of dark matter particles. By substituting (7) into (9), we obtain the following result:
$\left[-\dfrac{1}{mr}\dfrac{d^{2}}{dr^{2}}r-\dfrac{l(l+1)}{mr^{2}}-E+V(r)-i\Gamma\dfrac{N\delta(r)}{8\pi r^{2}}\right]G_{ba}^{(E,l)}(r,r^{\prime})=\dfrac{\delta(r-r^{\prime})}{r^{2}}\delta_{ba}$ (11)
In this analysis, we focus on the annihilation processes with orbital angular momentum $l=0$ (s-wave), which provide the leading-order contributions in the non-relativistic limit. Consequently, the total angular momentum $J$ is equal to the total spin angular momentum $S$. The leading-order potential terms exhibit spherical symmetry, resulting in the conservation of $J$ in annihilation processes. Furthermore, the conservation of total electric charge $Q$ is upheld. As a result, the Green function is expressed in the diagonal form for each eigenstate corresponding to $Q$ and $J$. To simplify the differential equation, we introduce $G_{ba}^{(E,l)}(r,r^{\prime})=\frac{g_{ba}(r,r^{\prime})}{rr^{\prime}}$.
$\left[-\dfrac{1}{mr}\dfrac{d^{2}}{dr^{2}}r-E+V(r)-i\Gamma\dfrac{N\delta(r)}{8\pi r^{2}}\right]g_{ba}^{(E,l)}(r,r^{\prime})=\delta(r-r^{\prime})\delta_{ba}$ (12)
The cross-section is determined by the imaginary part of the solution, and we derive the leading-order solution within the perturbative expansion of $\alpha$. The leading order solution can be written as
$g_{ba}(r,r^{\prime})|_{leadingorder}=\dfrac{iNm^{2}}{8\pi}\left[g_{>}(r^{\prime})\Gamma g_{>}^{T}(r)\right]_{ba}$ (13)
In the above expression, we introduce $g_{>ab}$ which is the solution of the following equation
$\left[-\dfrac{1}{m}\dfrac{d}{dr^{2}}-E+V(r)\right]g_{>ab}(r)=0$ (14)
with the boundary conditions
$g_{>ab}(r=0)=\delta_{ba}$ (15)
We impose different profiles depending on each potential. The short-range Yukawa potential diminishes as $r$ approaches infinity, causing the solution to exhibit an outgoing plane wave behavior. In the case of the long-range Coulomb potential, the asymptotic profile is determined by solving the Schrödinger equation in (14). In particular, we establish the following condition to govern the asymptotic behavior.
$g_{>ba}(r)\longrightarrow\ \left\{\begin{array}[]{lcc}d_{ba}(E_{a})e^{i\sqrt{mE_{b}}r}\hskip 28.45274ptshort-range~{}potential\\ \\ d_{ba}(E_{a})e^{i\sqrt{mE_{b}}r}e^{i\frac{m\alpha}{2\sqrt{mE_{b}}}\log{2\sqrt{mE_{b}}r}}\hskip 28.45274ptlarge-range~{}potential\\ \end{array}\right.$ (19)
Here, $E_{a}$ and $E_{b}$ denote the non-relativistic kinetic energy of $\phi_{a}$ and $\phi_{b}$, respectively. Given the constraints of numerical calculations, it is not feasible to let $r$ approach infinity. Therefore, we enforce these asymptotic behaviors at a sufficiently large value of $r=r_{f}$, ensuring that the numerical solutions become independent of $r_{f}$. The deviation from the asymptotic plane wave profile is quantified by $d_{ba}(E_{a})$, which is commonly referred to as the Sommerfeld Factor.
We numerically solve equation (14) to obtain the Sommerfeld factor as a function of the relative velocity for the initial state $v$. This effect becomes significant in the non-relativistic regime ($v\ll 1$), where the potential energy becomes comparable to the kinetic energy ($E\simeq\frac{mv^{2}}{4}$ ) in the Schrödinger equations. Conversely, the enhancement factors approach unity for $v\gg 1$. Consequently, we can recover the perturbative cross-section, as obtained by the Born-Oppenheimer approximation, by replacing the Sommerfeld factor with 1.
Figure 1: Sommerfeld enhancement effect
3.2 Annihilation Cross Section
The optical theorem enables us to establish a connection between $Im(G_{aa}^{E,l=0}(r,r^{\prime}))$ and the s-wave spin-averaged, velocity-weighted annihilation cross-section $\braket{\sigma_{a}v}$. This connection results in the following formula.
[…]
In the equation above, there is no summation taken over $a$, and $\sigma_{a}v$ represents the annihilation cross section of $\phi_{a}\sim V^{i}V^{j}$. We introduce a prefactor denoted as $c_{a}$, defined as follows:
[…]
If we substitute the leading order solution given in (13), we obtain
$\braket{\sigma_{a}v}=c_{a}\dfrac{m^{2}}{k^{2}}\lim_{E\longrightarrow\frac{k^{2}}{m^{2}}}\left(E-\dfrac{k^{2}}{m^{2}}\right)^{2}[\mathcal{A}(k)~{}\Gamma~{}\mathcal{A}^{\dagger}(k)]_{aa}$ (25)
where we define
[…]
The asymptotic behavior $r\longrightarrow\infty$ is important to evaluate $\mathcal{A}_{ba}(k)$. Substituting (19) into (26), we obtain
[…]
$=d_{ba}(E)\left(-\dfrac{k}{m}\right)\left(E-\frac{k^{2}}{m}\right)^{-1}$ (28)
We replace into (25), we express the annihilation cross section as below
$\braket{\sigma_{a}v}\simeq c_{a}\sum_{b,a}d_{ab}(E)\Gamma_{bc}d_{ac}^{\ast}(E)$ (29)
Specifically, we are interested in the velocity-weighted annihilation cross section for the partial annihilation processes $V^{0}V^{0}\longrightarrow XX^{\prime}$, where $X$ represents any Standard Model particle.
$\braket{\sigma_{a}v}_{XX^{\prime}}=c\sum_{\alpha,\beta}\sum_{J,J_{z}}(\Gamma_{XX^{\prime}}^{J,J_{z}})_{\alpha\beta}~{}d_{3\alpha}(E)d_{3\beta}^{\ast}(E)$ (30)
Here, $c$ is set to $2$ for initial states consisting of identical particles. We define $E$ as approximately $E\simeq\frac{mv_{rel}^{2}}{4}$ representing the non-relativistic kinetic energy of the V-particles, and $d_{\alpha\beta}(E)~{}(\alpha,\beta=1,2,3,4)$ as the Sommerfeld enhancement factor.
3.3 Non-Relativistic Limit
To ensure that the dark matter vector remains non-relativistic, we must carry out the following non-relativistic limit of the operators, as defined in [1].
$V_{\mu}^{0}(x)=\dfrac{1}{\sqrt{2m}}[v_{0}(\vec{x})\epsilon_{\mu}(x)e^{imt}+a_{0}^{\dagger}(\vec{x})\epsilon_{\mu}^{\ast}(x)e^{-imt}]\hskip 14.22636pt$ (31)
$(V_{\mu}^{0})^{\ast}(x)=\dfrac{1}{\sqrt{2m}}[a_{0}(\vec{x})\epsilon_{\mu}(x)e^{imt}+v_{0}^{\dagger}(\vec{x})\epsilon_{\mu}^{\ast}(x)e^{-imt}]\hskip 14.22636pt$ (32)
$V_{\mu}^{+}(x)=\dfrac{1}{\sqrt{2m}}[v_{+}(\vec{x})\epsilon_{\mu}(x)e^{imt}+v_{-}^{\dagger}(\vec{x})\epsilon_{\mu}^{\ast}(x)e^{-imt}]\hskip 14.22636pt$ (33)
[…]
Here, $v_{0},a_{0},v_{+},v_{-}$ represent the non-relativistic limits of the dark matter vector field. $m$ denotes the mass of the neutral component of the dark matter vector field, and $\epsilon_{\mu}$ signifies the polarization vector. These polarization vectors adhere to the transverse conditions and an orthonormal relation:
$p^{\mu}\epsilon_{\mu}=0;\hskip 28.45274pt\epsilon_{\mu}\epsilon_{\nu}=-\delta_{\mu\nu}$ (35)
3.4 Electroweak Matrix
3.4.1 Gauge Bosons-Dark Matter Interaction
We aim to investigate the annihilation of dark matter particles into standard model ones. Specifically, we consider the pairs ($v_{0}v_{0}$), for which such annihilation occurs. The most important contributions come from processes involving the exchange of $Z$ and $W^{\pm}$ bosons, having in their initial states ($v_{0}v_{0}$) , ($v_{0}a_{0}$), or ($v_{+}v_{-}$) pairs. Our approach involves deriving an effective action that accurately describes these pairs. For this purpose, we define the following vector:
$s(x,\vec{y})=\begin{pmatrix}\dfrac{v_{0}(x)v_{0}(x^{0},\vec{y})}{\sqrt{2}}\\ \dfrac{a_{0}(x)a_{0}(x^{0},\vec{y})}{\sqrt{2}}\\ v_{0}(x)a_{0}(x^{0},\vec{y})\\ v_{+}(x)v_{-}(x^{0},\vec{y}).\end{pmatrix}$ (40)
The disparity in the normalization of the first component in (40) arises from the fact that they pertain to states composed of two identical particles. The effective action is as follows:
$S_{eff,1}=$ $-\int\dfrac{d^{4}xd^{3}y}{8\pi|\vec{x}-\vec{y}|}(J_{A}^{0}(x)J_{A}^{0}(x^{0},\vec{y})$
$+J_{Z}^{0}(x)J_{Z}^{0}(x^{0},\vec{y})e^{-M_{Z}|\vec{x}-\vec{y}|}$
$+2J_{W_{+}}^{0}(x)J_{W_{-}}^{0}(x^{0},\vec{y})e^{-M_{W}|\vec{x}-\vec{y}|})$ (41)
The currents are calculated from the cubic interactions of the Lagrangian (2) between dark matter and gauge bosons. From these expressions, we employ the non-relativistic limit to obtain the electroweak matrix, which we will use to solve the Schrödinger equation (14).
In the non-relativistic limit, the first term of equation (3.4.1) characterizes the Coulomb interaction between non-relativistic particles, whereas the remaining terms depict the Yukawa potentials related to the exchange of $Z$ and $W$ bosons. Upon employing vector (40), the effective action transforms into the following expression:
$S_{eff,1_{NR}}=-\int d^{4}xd^{3}y~{}s(x,\vec{y})^{\dagger}V_{Gauge}(|\vec{x}-\vec{y}|)s(x,\vec{y})$ (42)
where $V_{Gauge}(|\vec{x}-\vec{y}|)$ is electroweak potential matrix (interaction gauge boson-vector dark matter), with $r=|\vec{x}-\vec{y}|$;
$V_{Gauge}(r)=-\dfrac{g^{2}}{8\pi r}\begin{pmatrix}4s_{W}^{2}e^{-M_{Z}r}&0&0&0\\ 0&4s_{W}^{2}e^{-M_{Z}r}&0&0\\ 0&0&16s_{W}^{2}e^{-M_{Z}r}&2\sqrt{2}e^{-M_{W}r}\\ 0&0&2\sqrt{2}e^{-M_{W}r}&8s_{W}^{2}e^{-M_{Z}r}-2(c_{W}-s_{W}t_{W}^{2})^{2}\end{pmatrix}$ (47)
In appendix A we show how to compute this matrix.
3.4.2 Higgs Boson-Dark Matter Interaction
We need to take into account the interaction between the Higgs Boson and the Dark Matter vector. Our primary focus lies in understanding the cubic interaction, as it is crucial for unraveling the nature of this interaction. For this purpose, we parametrize the Higgs doublet as folloes:
$\phi(x)=\begin{pmatrix}\phi^{+}(x)\\ \dfrac{v+h(x)}{\sqrt{2}}+i\phi^{0}(x)\end{pmatrix}$ (50)
where $\phi^{+}$ and $\phi^{0}$ are would-be Goldtone bosons, and $v$ is vaccum expectation value.
Now, when we substitute $V_{\mu}$, $V_{\mu}^{\dagger}$, $\phi$, and $\phi^{\dagger}$ into the Lagrangian (2), we can derive the electroweak matrix describing the interaction between the Higgs and the Dark Matter vector, which takes the form:
$V_{h-V}(r)=\dfrac{v^{2}}{32\pi m^{2}r}\begin{pmatrix}8\lambda_{4}^{2}a&0&0&0\\ 0&8\lambda_{4}^{2}a&0&0\\ 0&0&\dfrac{1}{2}(\lambda_{2}+\lambda_{3})^{2}a&(\lambda_{3}+2\lambda_{4})^{2}e^{-M_{W}r}\\ 0&0&(\lambda_{3}+2\lambda_{4})^{2}e^{-M_{W}r}&\lambda_{2}^{2}a\end{pmatrix}$ (55)
In Appendix A, we illustrate the methodology for computing this matrix.
3.5 Annihilation Matrix
3.5.1 Gauge Boson-Dark Matter Vector
When Dark Matter particles annihilate into Standard Model ones, it becomes essential to compute the annihilation matrix, specifically examining the interaction between Gauge Bosons and the Dark Matter. This requires the use of the quartic terms from Lagrangian (2). The effective action for these quartic terms, in the non-relativistic limit, can be expressed as follows:
$S_{eff,2NR}$ $=2i\int~{}d^{4}xd^{3}y~{}s(x,\vec{y})^{\dagger}\Gamma_{gauge}\delta^{3}(|\vec{x}-\vec{y}|)s(x,\vec{y})$ (56)
In Appendix B, the computation of these terms is detailed. Considering the different channels and total spin states, we obtain the following annihilation matrices:
$\Gamma_{WW}^{J=0}=\dfrac{2g^{4}}{64\sqrt{3}\pi m^{2}}\begin{pmatrix}2&0&0&0\\ 0&2&0&0\\ 0&0&1&1\\ 0&0&1&1\end{pmatrix}$ (61)
$\Gamma_{WW}^{J=2}=\dfrac{g^{4}}{64\pi m^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2&0&0&0\\ 0&2&0&0\\ 0&0&1&1\\ 0&0&1&1\end{pmatrix}$ (66)
$\Gamma_{AA}^{J=0}=\dfrac{e^{4}}{128\sqrt{3}\pi m^{2}}\begin{pmatrix}2a^{2}&0&0&0\\ 0&2a^{2}&0&0\\ 0&0&a^{2}&4a^{2}\\ 0&0&4a^{2}&4\end{pmatrix}$ (71)
$\Gamma_{AA}^{J=2}=\dfrac{e^{4}}{256\pi m^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2a^{2}&0&0&0\\ 0&2a^{2}&0&0\\ 0&0&a^{2}&4a^{2}\\ 0&0&4a^{2}&4\end{pmatrix}$ (76)
with $a=(1-\tan\theta_{W})$
$\Gamma_{ZZ}^{J=0}$ $=\dfrac{g^{4}}{128\sqrt{3}\pi m^{2}}\begin{pmatrix}2b^{2}&0&0&0\\ 0&2b^{2}&0&0\\ 0&0&2b^{2}&-4c\\ 0&0&-4c&4\sin^{4}\theta_{W}\end{pmatrix}$ (81)
$\Gamma_{ZZ}^{J=2}$ $=\dfrac{g^{4}}{256\pi m^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2b^{2}&0&0&0\\ 0&2b^{2}&0&0\\ 0&0&2b^{2}&-4c\\ 0&0&-4c&4\sin^{4}\theta_{W}\end{pmatrix}$ (86)
with $b=(\tan^{2}\theta_{W}\sin^{2}\theta_{W}-\sin\theta_{W}\cos\theta_{W})$ y $c=\sin^{2}\theta_{W}\tan\theta_{W}(\sin^{2}\theta_{W}-\cos^{2}\theta_{W})$
$\Gamma_{AZ}^{J=0}$ $=\dfrac{e^{2}g^{2}}{128\sqrt{3}\pi m^{2}}\begin{pmatrix}d^{2}&0&0&0\\ 0&d^{2}&0&0\\ 0&0&d^{2}&f\\ 0&0&f&\dfrac{\sin^{2}(2\theta_{W})}{\cos^{2}\theta_{W}}\end{pmatrix}$ (91)
$\Gamma_{AZ}^{J=2}$ $=\dfrac{e^{2}g^{2}}{256\pi m^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}d^{2}&0&0&0\\ 0&d^{2}&0&0\\ 0&0&d^{2}&f\\ 0&0&f&\dfrac{\sin^{2}(2\theta_{W})}{\cos^{2}\theta_{W}}\end{pmatrix}$ (96)
with $d=(\cos\theta_{W}-\sin\theta_{W})$ y $f=(\cos\theta_{W}-\sin\theta_{W})^{4}(1+\tan\theta_{W})^{2}$
3.5.2 Higgs Boson-Dark Matter Vector
We follow a similar procedure to take into accounts the Higgs Boson and the Dark Matter interactions. In this case, we require the quartic terms from Lagrangian (2), the quartic terms involving the new vector field and the Higgs boson doublet. The resulting annihilation matrix can be written as follows:
$\Gamma_{Higgs}=\dfrac{1}{32\pi m^{2}}\begin{pmatrix}\left(\dfrac{3}{2}+\sqrt{2}\right)^{2}\lambda_{4}^{2}&0&0&0\\ 0&\dfrac{9}{4}\lambda_{4}^{2}&0&0\\ 0&0&\dfrac{1}{4}(7\lambda_{2}^{2}+6\lambda_{2}\lambda_{3}+5\lambda_{3}^{2})&\dfrac{1}{2}(\lambda_{3}+2\lambda_{4})^{2}\\ 0&0&\dfrac{1}{2}(\lambda_{3}+2\lambda_{4})^{2}&\dfrac{3}{4}\lambda_{2}^{2}+(\lambda_{2}+\lambda_{3})^{2}\end{pmatrix}$ (101)
3.6 Results
Now, we have all the needed ingredients to compute the thermal averaged cross-section $\braket{\sigma v}$ using equation (30). When we expand over all the annihilation channels, $\braket{\sigma v}$ takes the form:
$\braket{\sigma v}=\braket{\sigma v}_{W^{+}W^{-}}+\braket{\sigma v}_{ZZ}+\braket{\sigma v}_{\gamma\gamma}+\braket{\sigma v}_{Z\gamma}+\braket{\sigma v}_{hh}.$ (102)
Our results are shown in figure 2 . There are several noteworthy characteristics of the Sommerfeld enhancement that merit discussion. Firstly, it is a velocity-dependent phenomenon. For instance, no structure is observed for a relative velocity of $v=0.3~{}c$ (dashed lines in figure 2), which is representative of the early Universe, when the relic density was formed. In contrast, for smaller velocities, like the one expected in the galaxy ($v=10^{-3}c$), significant oscillations and resonances appear. For masses that exhibit such resonances (roughly around $6$ TeV), the Sommerfeld factor experiences a significant boost of about 5 order of magnitude, for all the values of portal coupling constant $\lambda$.
Figure 2: Annihilation cross-section (fundamental representation spin-1 of $SU(2)_{L}$) for the process $v_{0}a_{0}\longrightarrow SMSM$ with the Sommerfeld Enhancement factor, considering velocities $v=10^{-3}$ (galactic Halo, shown as continuous lines, green, blue, black and gray lines ) and $v=0.3$ (Relic density, indicated by dashed lines), and different values of the Higgs portal coupling constant $\lambda$. The green line indicates that it is the spin-1 triplet representation of $SU(2)_{L}$, while the red line indicates that it is the spin-1/2 doublet representation of $SU(2)_{L}$.
4 Photons Flux from Annihilation neutral Dark Matter particles
The $\gamma$-ray flux resulting from the annihilation of neutral Dark Matter particles, as described in [14], in a region of the sky with a solid angle $\Omega$ is given by:
$\dfrac{d\Phi}{dE}=\dfrac{1}{8\pi}J(\Omega)\dfrac{\braket{\sigma v}_{\gamma\gamma}}{m^{2}}\left(\dfrac{dN}{dE}\right)_{\gamma\gamma}$ (103)
Here, $J(\Omega)$ represents the astrophysical $J$-factor, which measures the relevance and interest of the astrophysical target for Dark Matter annihilation.
[…]
The $J$-factors for dwarf galaxies are approximately $J\approx 10^{19}$ $GeV^{2}/cm^{5}$, for the Andromeda galaxy, $J\approx 10^{20}$ $GeV^{2}/cm^{5}$, and for the Milky Way’s Galactic Center, $J\approx 10^{25}$ $GeV^{2}/cm^{5}$. If the solid angle changes by one degree, the $J$-factor undergoes an approximate change of $J\approx 10^{24}$ $GeV^{2}/cm^{5}$.
The distribution, denoted as $\left(\frac{dN}{dE}\right)_{\gamma\gamma}$, characterizes the number of photons with energy $E_{\gamma\gamma}=m$, where $m$ represents the mass of the WIMP. The distribution is defined as follows:
$\left(\dfrac{dN}{dE}\right)_{\gamma\gamma}=2\times\delta(E-E_{\gamma\gamma})$ (105)
while $\braket{\sigma v}_{\gamma\gamma}$ represents the annihilation cross-section rate with the mediator, which is dependent on the branching ratio, and is defined as $\braket{\sigma v}_{\gamma\gamma}=BR_{\gamma\gamma}\times\braket{\sigma v}_{XX^{\prime}}$. Here, $\braket{\sigma v}_{XX^{\prime}}$ corresponds to the total cross-section for Dark Matter annihilation, as shown in equation (30).
$BR_{\gamma\gamma}=\dfrac{\Gamma_{\gamma\gamma}}{\sum_{V}\Gamma_{VV^{\prime}}}$ (106)
where $V$ is the gauge and Higgs bosons. Branching ratio is the quotient between annihilation matrix of photons by add total annihilation matrix.
4.1 Results
Using the expressions above, we calculate the photon flux resulting from the Annihilation Cross Section (30), which serves as a method for the Indirect Detection of Dark Matter. This entails the integration of equation (103) using the distribution of the number of photons as defined in (105). Figure 3 depict the photon flux from VDM annihilation per year. We can see, as expected, that as the Higgs portal increases in size, the photon flux also increases.Our constructions are not constrained by Fermi-LAT, as our region of interest is above $100$ GeV which is not tested by Fermi-LAT; therefore, in the graphs, we compare with the projected CTA sensitivity [16]. We can see that for low vlues of the Higgs-DM coupling CTA can test only the region around the big resonance while for larger values CTA will be able to observe or rule out all our region of interest.
Figure 3: Photons Flux from Annihilation of neutral Dark Matter particles (Yellow, Dark Orange, Purple and Dark Pink lines correspond to fundamental representation of $SU(2)_{L}$ spin-1), black line is the projected CTA photon flux sensitivity, The green line corresponds to the photon flux of the spin-1 $SU(2)_{L}$ triplet representation, while the red line corresponds to the $SU(2)_{L}$ doublet representation for spin-1/2.
5 Summary
In this paper, we have studied the Sommerfeld enhancement in a model where the DM candidate is a neutral component of a new massive vector field in the fundamnetal representation of $SU(2)_{L}$. We found that big resonance in $\sigma v$ for DM masses in the rage of 5 TeV to 10 TeV, in the low velocities regime. We used this result to predict the photon flux produced by DM annihilation at the center of the galaxy. This prediction can be tested in the near future by CTA.
Acknowledgments
This work was partially found by Fondecyt grant 1230110, ANID PIA/APOYO AFB220004, and ANID—The Millennium Science Initiative Program ICN2019_044. We are very thankful to Motoko Fujiwara for discussions and suggestions that contributed significantly to improve the quality of this paper.
Appendix A Electroweak Matrix
A.1 Gauge-Dark Matter
We are required to calculate the electroweak matrix using the effective action derived from the Lagrangian (2), resulting in the following expression:
$S_{eff,1}=-\frac{1}{2}\int d^{4}xd^{4}y~{}J^{\mu\dagger}(y)\Delta_{0~{}\mu\nu}^{-1}(x,y)J^{\nu}(x)$ (107)
Now, we proceed to compute $S_{eff,1}$ as given in equation (107) in the non-relativistic limit, commencing with:
$\Delta_{0~{}\mu\nu}^{-1}(x,y)=-\int\dfrac{d^{4}q}{(2\pi)^{4}}e^{iq(x-y)}\begin{pmatrix}\frac{1}{q^{2}+i\epsilon}&0&0&0\\ 0&\frac{1}{q^{2}-M_{Z}^{2}+i\epsilon}&0&0\\ 0&0&\frac{1}{q^{2}-M_{W}^{2}+i\epsilon}&0\\ 0&0&0&\frac{1}{q^{2}-M_{W}^{2}+i\epsilon}\end{pmatrix}g_{\mu\nu}$ (112)
Next, we substitute the gauge currents and (112) into (107), resulting in:
$S_{eff,1}=\int\dfrac{d^{4}xd^{4}yd^{4}q}{2(2\pi)^{4}}e^{iq(x-y)}\left(\dfrac{J^{\mu}_{A}(x)J_{A~{}\mu}(y)}{q^{2}+i\epsilon}+\dfrac{J^{\mu}_{Z}(x)J_{Z~{}\mu}(y)}{q^{2}-M_{Z}^{2}+i\epsilon}+\dfrac{2J^{\mu}_{W^{+}}(x)J_{W^{-}~{}\mu}(y)}{q^{2}-M_{W}^{2}+i\epsilon}\right)$ (113)
Now, we need to determine the non-relativistic limit of (113). However, we have the relativistic gauge currents expressed as $J_{V}=(J_{V}^{0},\vec{J_{V}}),$ where $V$ represents any gauge boson. With this representation, each term of (113) can be expressed as:
$\dfrac{J^{\mu}_{V}(x)J_{V~{}\mu}(y)}{q^{2}-M_{V}^{2}+i\epsilon}=-\dfrac{J^{0}_{V}(x)J_{V}^{0}(y)}{\vec{q}^{2}+i\epsilon}+\dfrac{J^{0}_{V}(x)J_{V}^{0}(y)q^{0^{2}}}{(\vec{q}^{2}+i\epsilon)(q^{2}-M_{V}^{2}+i\epsilon)}+\dfrac{\vec{J}^{\mu}_{V}(x)\vec{J}_{V\mu}(y)}{q^{2}-M_{V}^{2}+i\epsilon}$ (114)
The third term in equation (114) involves magnetic interactions between currents. To account for non-relativistic currents, we require that $|\vec{J_{V}}|\ll J_{V}^{0}$, and as a result, this term can be neglected. The second term is associated with the retarded effects on the interactions between currents. Therefore, we must perform integration by parts on this term, leading to:
$\int$ $d^{4}xd^{4}yd^{4}q~{}e^{iq(x-y)}\dfrac{J_{V}^{0}(x)J_{V}^{0}(y)q^{0^{2}}}{(\vec{q}^{2}+M_{V}^{2})(q^{2}-M_{V}^{2}+i\epsilon)}$
=$\int$= ∫ $d^{4}xd^{4}yd^{4}q~{}e^{iq(x-y)}\dfrac{\dot{J}_{V}^{0}(x)\dot{J}_{V}^{0}(y)}{(\vec{q}^{2}+M_{V}^{2})(q^{2}-M_{V}^{2}+i\epsilon)}$ (115)
Using the condition $\mid\frac{1}{v_{i}}\frac{dv_{i}}{dt}\mid\ll M_{V_{0}},$ we can deduce that $|\dot{J}{V}^{0}|\ll J{V}^{0}$, and the retarded effects render this term negligible in the non-relativistic limit. Consequently, we need to integrate the first term.
$\int\dfrac{d^{4}q}{(2\pi)^{4}}\dfrac{e^{iq(x-y)}}{\vec{q}^{2}+M_{V}^{2}}=\delta(x^{0}-y^{0})\int\dfrac{d^{3}q}{(2\pi)^{3}}\dfrac{e^{iq(x-y)}}{\vec{q}^{2}+M_{V}^{2}}=\delta(x^{0}-y^{0})\dfrac{e^{-M_{V}|\vec{x}-\vec{y}|}}{4\pi|\vec{x}-\vec{y}|}$ (116)
We substitute equation (116) into equation (113), resulting in the following effective action in the non-relativistic limit:
$S_{eff,1}=-\int\dfrac{d^{4}xd^{3}y}{8\pi|\vec{x}-\vec{y}|}($ $J_{A}^{0}(x)J_{A}^{0}(x^{0},\vec{y})+J_{Z}^{0}(x)J_{Z}^{0}(x^{0},\vec{y})e^{-M_{Z}|\vec{x}-\vec{y}|}$
$+2J_{W^{+}}^{0}(x)J_{W^{-}}^{0}(x^{0},\vec{y})e^{-M_{W}|\vec{x}-\vec{y}|})$ (117)
The first term in equation (A.1) accounts for the Coulomb potential among non-relativistic particles, whereas the remaining terms describe Yukawa potential associated with the exchange of $W$ and $Z$ bosons.
It is essential to consider only pairs in the form of vectors, as indicated in (40), while other particle combinations are prohibited. Consequently, the currents in the non-relativistic limit are as follows:
$J_{A}^{0}(x)J_{A}^{0}(x^{0},\vec{y})=2e^{2}s_{W}^{2}(c_{W}-s_{W}t_{W}^{2})$ $(-[v_{+}^{\dagger}(x)v_{-}^{\dagger}(x^{0},\vec{y})v_{+}(x)v_{-}(x^{0},\vec{y})]$ (118)
$J_{Z}^{0}(x)J_{Z}^{0}(x^{0},\vec{y})$ $=-g^{2}s_{W}^{2}\theta_{W}[v_{+}^{\dagger}(x)v_{-}^{\dagger}(x^{0},\vec{y})v_{+}(x)v_{-}(x^{0},\vec{y})$
$+\left(\dfrac{v_{0}^{\dagger}(x)v_{0}^{\dagger}(x^{0},\vec{y})}{\sqrt{2}}\dfrac{v_{0}(x)v_{0}(x^{0},\vec{y})}{\sqrt{2}}\right)$
$+\left(\dfrac{a_{0}^{\dagger}(x)a_{0}^{\dagger}(x^{0},\vec{y})}{\sqrt{2}}\dfrac{a_{0}(x)a_{0}(x^{0},\vec{y})}{\sqrt{2}}\right)$
$-\left(v_{0}^{\dagger}(x)a_{0}^{\dagger}(x^{0},\vec{y})v_{0}(x)a_{0}(x^{0},\vec{y})\right)]$ (119)
$J_{W_{+}}^{0}(x)J_{W_{-}}^{0}(x^{0},\vec{y})=-\dfrac{g}{2}\left[v_{0}^{\dagger}(x)a_{0}^{\dagger}(x^{0},\vec{y})v_{+}(x)v_{-}(x^{0},\vec{y})-v_{-}^{\dagger}(x)v_{+}^{\dagger}(x^{0},\vec{y})a_{0}(x)v_{0}(x^{0},\vec{y})\right]$ (120)
where $s_{W}=\sin\theta_{W}$, $c_{W}=\cos\theta_{W}$ and $t_{W}=\tan\theta_{W}$. Here the terms that are omitted are the ones that do not lead to pairs $v_{0}v_{0}$, $a_{0}a_{0}$, $v_{0}a_{0}$ and $v_{+}v_{-}$. By plugging these currents in equation (A.1), we obtain:
$S_{eff,1_{NR}}=-\int d^{4}xd^{3}y~{}s(x,\vec{y})^{\dagger}V_{Gauge}(|\vec{x}-\vec{y}|)s(x,\vec{y})$ (121)
where $V_{Gauge}(|\vec{x}-\vec{y}|)$ is the electroweak matrix, with $r=|\vec{x}-\vec{y}|$, therefore:
$V_{Gauge}(r)=-\dfrac{g^{2}}{8\pi r}\begin{pmatrix}4s_{W}^{2}e^{-M_{Z}r}&0&0&0\\ 0&4s_{W}^{2}e^{-M_{Z}r}&0&0\\ 0&0&16s_{W}^{2}e^{-M_{Z}r}&2\sqrt{2}e^{-M_{W}r}\\ 0&0&2\sqrt{2}e^{-M_{W}r}&8s_{W}^{2}e^{-M_{Z}r}-2(c_{W}-s_{W}t_{W}^{2})^{2}\end{pmatrix}$ (126)
A.2 Higgs Boson-Dark Matter
We are now employing the Feynman Gauge. The masses of the Goldstone bosons, namely $\phi^{0}$, $\phi^{0\ast}$, and ϕ±superscriptitalic-$\phi^{\pm}$, originating from the doublet (50), are $M_{Z}$ and $M_{W}$. The Higgs particle is significantly lighter than Dark Matter particles. The interacting Lagrangian (2) between the Higgs Boson and the Dark Matter Vector is as follows:
$\mathcal{L}_{h-V}=-\lambda_{2}(\phi^{\dagger}\phi)(V_{\mu}^{\dagger}V^{\mu})-\lambda_{3}(\phi^{\dagger}V_{\mu})(V^{\mu\dagger}\phi)-\dfrac{\lambda_{4}}{2}\left[(\phi^{\dagger}V_{\mu})(\phi^{\dagger}V^{\mu})+(V^{\mu\dagger}\phi)(V_{\mu}^{\dagger}\phi)\right]$ (127)
In order to construct pairs such as $v_{0}v_{0},v_{0}a_{0},a_{0}a_{0}$, and $v_{+}v_{-}$, we need to substitute equations (5) and (50) into (127). Consequently, we obtain:
$V_{vector}(r)=\dfrac{v^{2}}{32\pi m^{2}r}\begin{pmatrix}8\lambda_{4}^{2}a&0&0&0\\ 0&8\lambda_{4}^{2}a&0&0\\ 0&0&\dfrac{1}{2}(\lambda_{2}+\lambda_{3})^{2}a&(\lambda_{3}+2\lambda_{4})^{2}e^{-M_{W}r}\\ 0&0&(\lambda_{3}+2\lambda_{4})^{2}e^{-M_{W}r}&\lambda_{2}^{2}a\end{pmatrix}$ (132)
this is electroweak matrix (Higgs Boson-Dark Matter interaction)
Appendix B Annihilation Matrix
B.1 Gauge-Dark Matter Interaction
The non-relativistic vector dark matter particles have the potential to annihilate into relativistic particles, specifically gauge bosons of $SU(2){L}$. In the framework of effective quantum field theory, these annihilations are depicted by the inclusion of an imaginary potential or an absorptive term in the action, which accounts for the disappearance of pairs of non-relativistic particles. These absorptive terms originate from $S{eff,2}$ as outlined in equation (56). To provide an interpretation, we can visualize $S_{eff,2}$ in terms of Feynman diagrams. In this representation, the matrix $\Delta_{0}^{-1}$ characterizes the propagators of gauge fields, and each $\mathcal{Q}$ represents a vertex featuring two gauge fields and two vector dark matter fields, as described by Chowdhury et al. [7].
Figure 4: Feynman diagram of quartic terms
Then, the effective action of absorptive terms is:
$S_{eff,2}$ $=-iTr\log\left(\Delta_{0}^{-1}\mathcal{Q}\Delta_{0}^{-1}\mathcal{Q}\right)$
$=-4i\int d^{4}xd^{4}yd^{4}zd^{4}w\dfrac{d^{4}q}{(2\pi)^{4}}\dfrac{d^{4}k}{(2\pi)^{4}}\dfrac{e^{iq(x-y)}}{q^{2}+i\epsilon}\dfrac{e^{ik(w-z)}}{k^{2}+i\epsilon}Tr\left(\mathcal{Q}(y,z)\mathcal{Q}(w,x)\right)$ (133)
The factor of 4 corresponds to the trace of gauge fields’ indices, and our focus is on the imaginary part. To proceed, we consider the gauge bosons in the loop to be on-shell, as follows:
$\dfrac{1}{q^{2}+i\epsilon}\rightarrow-2\pi i\delta(q^{2})$ (134)
Therefore, we perform an integration over the timelike component of the 4-momentum, resulting in the transformation $d^{4}q\delta(q^{2})\rightarrow\frac{d^{3}q}{2|\vec{q}|}$ when the gauge particles are on-shell. The integration of 4-momentum is then expressed as:
$\int\dfrac{d^{4}q}{(2\pi)^{4}}\dfrac{d^{4}k}{(2\pi)^{4}}\dfrac{e^{iq(x-y)}}{q^{2}+i\epsilon}\dfrac{e^{ik(w-z)}}{k^{2}+i\epsilon}=-\dfrac{i}{8\pi}\delta^{4}(x-y)$ (135)
Then, replace into equation (B.1), we can obtain:
$S_{eff,2}=\frac{i}{2\pi}\int d^{4}xTr\begin{pmatrix}Q_{AA}(x)&Q_{AZ}(x)&Q_{AW}(x)&Q_{AW}^{\ast}(x)\\ Q_{AZ}(x)&Q_{ZZ}(x)&Q_{ZW}(x)&Q_{ZW}^{\ast}(x)\\ Q_{AW}^{\ast}(x)&Q_{ZW}^{\ast}(x)&Q_{WW}(x)&0\\ Q_{AW}^{\ast}(x)&Q_{ZW}^{\ast}(x)&0&Q_{WW}^{\ast}(x)\\ \end{pmatrix}^{2}$ (140)
We need only pairs like $V_{0}V_{0},V_{0}^{\ast}V_{0}^{\ast},V_{0}V_{0}^{\ast}$ and $V_{+}V_{-}$, then
$S_{eff,2}$ $=\dfrac{i}{2\pi}\int~{}d^{4}x\left(Q_{AA}(x)^{2}+Q_{ZZ}(x)^{2}+2Q_{WW}(x)^{2}+2Q_{AZ}(x)^{2}\right)$
$=-\int~{}d^{4}x\dfrac{g^{2}}{8}[2W_{\mu}^{+}V_{\nu}^{-}V^{+~{}\nu}W^{-~{}\mu}+2W_{\mu}^{-}(V_{\nu}^{0})^{\ast}V^{0~{}\nu}W^{+~{}\mu}$
$+A_{\mu}\left\{\sin^{2}\theta_{W}(1-\tan\theta_{W})(V_{\nu}^{0})^{\ast}V^{0~{}\nu}+2\sin^{2}\theta_{W}V_{\nu}^{-}V^{+~{}\nu}\right\}A^{\mu}$
$+Z_{\mu}\left\{(\tan^{2}\theta_{W}\sin^{2}\theta_{W}-\sin\theta_{W}\cos\theta_{W})(V_{\nu}^{0})^{\ast}V^{0~{}\nu}+2\sin^{2}\theta_{W}V_{\nu}^{-}V^{+~{}\nu}\right\}Z^{\mu}$
$+A_{\mu}((\sin\theta_{W}(2\cos\theta_{W}-\sin\theta_{W}(1+\tan\theta_{W}))(V_{\nu}^{0})^{\ast}V^{0~{}\nu}$
$+\cos(2\theta_{W})\tan\theta_{W}V_{\nu}^{-}V^{+~{}\nu})Z^{\mu}$
$+Z_{\mu}(\sin\theta_{W}(\cos\theta_{W}-\sin\theta_{W})(V_{\nu}^{0})^{\ast}V^{0~{}\nu}$
$+\tan\theta_{W}(\cos^{2}\theta_{W}-\sin^{2}\theta_{W})V_{\nu}^{-}V^{+~{}\nu})A^{\mu}](1+1+1+1)$
$=-\dfrac{g^{2}}{2}[2W_{\mu}^{+}V_{\nu}^{-}V^{+~{}\nu}W^{+~{}\mu}+2W_{\mu}^{-}(V_{\nu}^{0})^{\ast}V^{0~{}\nu}W^{+~{}\mu}$
$+A_{\mu}\left\{\sin^{2}\theta_{W}(1-\tan\theta_{W})(V_{\nu}^{0})^{\ast}V^{0~{}\nu}+2\sin^{2}\theta_{W}V_{\nu}^{-}V^{+~{}\nu}\right\}A^{\mu}$
$+Z_{\mu}\left\{(\tan^{2}\theta_{W}\sin^{2}\theta_{W}-\sin\theta_{W}\cos\theta_{W})(V_{\nu}^{0})^{\ast}V^{0~{}\nu}+2\sin^{2}\theta_{W}V_{\nu}^{-}V^{+~{}\nu}\right\}Z^{\mu}$
$+A_{\mu}((\sin\theta_{W}(2\cos\theta_{W}-\sin\theta_{W}(1+\tan\theta_{W}))(V_{\nu}^{0})^{\ast}V^{0~{}\nu}$
$+\cos(2\theta_{W})\tan\theta_{W}V_{\nu}^{-}V^{+~{}\nu})Z^{\mu}$
$+Z_{\mu}(\sin\theta_{W}(\cos\theta_{W}-\sin\theta_{W})(V_{\nu}^{0})^{\ast}V^{0~{}\nu}$
$+\tan\theta_{W}(\cos^{2}\theta_{W}-\sin^{2}\theta_{W})V_{\nu}^{-}V^{+~{}\nu})A^{\mu}]$ (141)
In terms of s’vectors, the effective action is
$S_{eff,2}=2i\int d^{4}xd^{3}y~{}s(x,\vec{y})^{\dagger}\Gamma_{gauge}\delta^{3}(|x-y|)s(x,\vec{y})$ (142)
The Q’s terms are:
$Q_{WW}^{2}(x)$ $=\dfrac{g^{4}}{16}(V_{\nu}^{-}V^{+~{}\nu}+(V_{\nu}^{0})^{\ast}V^{0~{}\nu})^{2}$ (143)
$=\dfrac{g^{4}}{16}(V_{i}^{-}V^{+~{}j}V_{i}^{-}V^{+~{}j}+(V_{i}^{0})^{\ast}V^{0~{}j}(V_{i}^{0})^{\ast}V^{0~{}j}$
$+2(V_{i}^{0})^{\ast}V^{0~{}j}(V_{i}^{-})^{\ast}V^{+~{}j})$
$Q_{AA}^{2}(x)=\dfrac{e^{4}}{64}\left[(1-\tan\theta_{W})^{2}((V_{\nu}^{0})^{\ast}V^{0~{}\nu})^{2}+4(V_{\nu}^{-}V^{+~{}\nu})^{2}+4(1-\tan\theta_{W})(V_{\nu}^{0})^{\ast}V^{0~{}\nu}V_{\nu}^{-}V^{+~{}\nu}\right]$ (144)
$Q_{ZZ}^{2}(x)=\dfrac{g^{4}}{64}$ $[(\tan^{2}\theta_{W}\sin^{2}\theta_{W}-\sin\theta_{W}\cos\theta_{W})^{2}((V_{\nu}^{0})^{\ast}V^{0~{}\nu})^{2}+4\sin^{4}\theta_{W}(V_{\nu}^{-}V^{+~{}\nu})^{2}$
$-4\sin^{2}\theta_{W}(\tan^{2}\theta_{W}\sin^{2}\theta_{W}-\sin\theta_{W}\cos\theta_{W})^{2}(V_{\nu}^{0})^{\ast}V^{0~{}\nu}V_{\nu}^{-}V^{+~{}\nu}]$ (145)
$Q_{AZ}^{2}(x)$ $=\dfrac{e^{2}g^{2}}{64}[(\cos\theta_{W}-\sin\theta_{W})^{2}((V_{\nu}^{0})^{\ast}V^{0~{}\nu})^{2}+\dfrac{\cos^{2}(2\theta_{W})}{\cos^{2}\theta_{W}}{\cos\theta_{W}}(V_{\nu}^{-}V^{+~{}\nu})^{2}$
$+(\cos\theta_{W}-\sin\theta_{W})^{4}(1+\tan\theta_{W})^{2}(V_{\nu}^{0})^{\ast}V^{0~{}\nu}V_{\nu}^{-}V^{+~{}\nu}]$ (146)
However, we need to establish a relationship between polarization vectors and angular momentum, as discussed in [1], where:
$\epsilon_{i}(p_{-})\epsilon_{j}(p_{+})\longrightarrow S_{ij}^{J,J_{z}}$ (147)
with add angular momentun matrices
$S^{0,0}=-\dfrac{1}{\sqrt{3}}\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix}$ (151)
$S^{1,1}=\dfrac{1}{2}\begin{pmatrix}0&0&-1\\ 0&0&-i\\ 1&i&0\end{pmatrix}\hskip 28.45274ptS^{1,0}=\dfrac{1}{\sqrt{2}}\begin{pmatrix}0&i&0\\ -i&0&0\\ 0&0&0\end{pmatrix}\hskip 28.45274ptS^{1,-1}=\dfrac{1}{2}\begin{pmatrix}0&0&-1\\ 0&0&i\\ 1&-i&0\end{pmatrix}$ (161)
$S^{2,2}=\dfrac{1}{2}\begin{pmatrix}1&i&0\\ i&-1&0\\ 0&0&0\end{pmatrix}\hskip 28.45274ptS^{2,1}=\dfrac{1}{2}\begin{pmatrix}0&0&-1\\ 0&0&-i\\ -1&-i&0\end{pmatrix}\hskip 28.45274ptS^{2,0}=\dfrac{1}{\sqrt{6}}\begin{pmatrix}-1&0&0\\ 0&-1&0\\ 0&0&2\end{pmatrix}$
$S^{2,-1}=\dfrac{1}{2}\begin{pmatrix}0&0&1\\ 0&0&-i\\ 1&-i&0\end{pmatrix}\hskip 28.45274ptS^{2,-2}=\dfrac{1}{2}\begin{pmatrix}1&-i&0\\ -i&-1&0\\ 0&0&0\end{pmatrix}$ (168)
We are using polarization vector for photons as:
$\epsilon_{i}^{\pm}(k_{1})=\dfrac{1}{\sqrt{2}}\begin{pmatrix}\mp\cos\theta\\ -i\\ \pm\sin\theta\end{pmatrix}\hskip 28.45274pt\epsilon_{i}^{\pm}(k_{2})=\dfrac{1}{\sqrt{2}}\begin{pmatrix}\pm\cos\theta\\ -i\\ \mp\sin\theta\end{pmatrix}$ (175)
The annilihation matrices (Gauge-Dark Matter) are:
$\Gamma_{WW}^{J=0}=\dfrac{2g^{4}}{64\sqrt{3}\pi m^{2}}\begin{pmatrix}2&0&0&0\\ 0&2&0&0\\ 0&0&1&1\\ 0&0&1&1\end{pmatrix}$ (180)
$\Gamma_{WW}^{J=2}=\dfrac{g^{4}}{64\pi m^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2&0&0&0\\ 0&2&0&0\\ 0&0&1&1\\ 0&0&1&1\end{pmatrix}$ (185)
$\Gamma_{AA}^{J=0}=\dfrac{e^{4}}{128\sqrt{3}\pi m^{2}}\begin{pmatrix}2a^{2}&0&0&0\\ 0&2a^{2}&0&0\\ 0&0&a^{2}&4a^{2}\\ 0&0&4a^{2}&4\end{pmatrix}$ (190)
$\Gamma_{AA}^{J=2}=\dfrac{e^{4}}{256\pi m^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2a^{2}&0&0&0\\ 0&2a^{2}&0&0\\ 0&0&a^{2}&4a^{2}\\ 0&0&4a^{2}&4\end{pmatrix}$ (195)
with $a=(1-\tan\theta_{W})$
$\Gamma_{ZZ}^{J=0}$ $=\dfrac{g^{4}}{128\sqrt{3}\pi m^{2}}\begin{pmatrix}2b^{2}&0&0&0\\ 0&2b^{2}&0&0\\ 0&0&2b^{2}&-4c\\ 0&0&-4c&4\sin^{4}\theta_{W}\end{pmatrix}$ (200)
$\Gamma_{ZZ}^{J=2}$ $=\dfrac{g^{4}}{256\pi m^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2b^{2}&0&0&0\\ 0&2b^{2}&0&0\\ 0&0&2b^{2}&-4c\\ 0&0&-4c&4\sin^{4}\theta_{W}\end{pmatrix}$ (205)
with $b=(\tan^{2}\theta_{W}\sin^{2}\theta_{W}-\sin\theta_{W}\cos\theta_{W})$ y $c=\sin^{2}\theta_{W}\tan\theta_{W}(\sin^{2}\theta_{W}-\cos^{2}\theta_{W})$
$\Gamma_{AZ}^{J=0}$ $=\dfrac{e^{2}g^{2}}{128\sqrt{3}\pi m^{2}}\begin{pmatrix}d^{2}&0&0&0\\ 0&d^{2}&0&0\\ 0&0&d^{2}&f\\ 0&0&f&\dfrac{\sin^{2}(2\theta_{W})}{\cos^{2}\theta_{W}}\end{pmatrix}$ (210)
$\Gamma_{AZ}^{J=2}$ $=\dfrac{e^{2}g^{2}}{256\pi m^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}d^{2}&0&0&0\\ 0&d^{2}&0&0\\ 0&0&d^{2}&f\\ 0&0&f&\dfrac{\sin^{2}(2\theta_{W})}{\cos^{2}\theta_{W}}\end{pmatrix}$ (215)
with $d=(\cos\theta_{W}-\sin\theta_{W})$ y $f=(\cos\theta_{W}-\sin\theta_{W})^{4}(1+\tan\theta_{W})^{2}$
B.2 Higgs-Dark Matter Interaction
From the Lagrangian (127), we can extract the quartic terms in the effective action. These terms provide the annihilation matrix for the interaction between the Higgs boson and the Dark Matter Vector. We need to operate on the vector in $SU(2)_{L}$ (5) and the Higgs boson doublet (50). Consequently,
$\mathcal{L}_{Quartic~{}h-V}$ $=-\lambda_{2}[\left(\phi^{-}\phi^{+}+\dfrac{h^{2}}{2}-i\dfrac{(\phi^{0\ast}h+h^{\ast}\phi^{0})}{\sqrt{2}}-(\phi^{0})^{2}\right)V_{\mu}^{-}V^{+\mu}$ (216)
$+\left(\phi^{-}\phi^{+}+\dfrac{h^{2}}{2}-i\dfrac{(\phi^{0\ast}h+h^{\ast}\phi^{0})}{\sqrt{2}}-(\phi^{0})^{2}\right)V_{\mu}^{0\ast}V^{0\mu}]$
$-\lambda_{3}[\phi^{-}\phi^{+}V_{\mu}^{+}V^{-\mu}+\dfrac{1}{\sqrt{2}}(i\phi^{0}+h)\phi^{-}V_{\mu}^{+}V^{0\mu\ast}+\dfrac{1}{\sqrt{2}}(-i\phi^{0\ast}+h^{\ast})\phi^{+}V_{\mu}^{0}V^{-\mu}$
$+\left(\dfrac{h^{2}}{\sqrt{2}}-(\phi^{0})^{2}\right)V_{\mu}^{0}V^{0\mu\ast}]$
$+\left(\dfrac{h^{2}}{\sqrt{2}}-(\phi^{0})^{2}\right)V_{\mu}^{0}V^{0\mu\ast}]$
$-\lambda_{4}[\left(-i\phi^{0\ast}+\dfrac{h^{\ast}}{\sqrt{2}}\right)\phi^{-}V^{\mu+}V^{0\mu}+\left(-i\phi^{0\ast}+\dfrac{h^{\ast}}{\sqrt{2}}\right)\phi^{-}V^{\mu 0}V^{+\mu}$
$+\left(\dfrac{h^{\ast}}{2}-i\sqrt{2}h^{\ast}\phi^{0\ast}+(\phi^{0\ast})^{2}\right)V_{\mu}^{0}V^{0\mu}+\left(i\phi^{0}+\dfrac{h}{\sqrt{2}}\right)\phi^{+}V^{-\mu}V_{\mu}^{0\ast}$
$+\left(i\phi^{0}+\dfrac{h}{\sqrt{2}}\right)\phi^{+}V^{0\mu\ast}V_{\mu}^{-}+\left(-(\phi^{0})^{2}+\dfrac{h^{2}}{2}\right)V^{0\mu\ast}V_{\mu}^{0\ast}]$
The permissible terms are $V_{0}V_{0},V_{0}V_{0}^{\ast},V_{0}^{\ast}V_{0}^{\ast}$, and $V_{+}V_{-}$, as we require neutral terms. The resulting annihilation matrix is as follows:
$\Gamma_{vector}=\dfrac{1}{32\pi m^{2}}\begin{pmatrix}\left(\dfrac{3}{2}+\sqrt{2}\right)^{2}\lambda_{4}^{2}&0&0&0\\ 0&\dfrac{9}{4}\lambda_{4}^{2}&0&0\\ 0&0&\dfrac{1}{4}(7\lambda_{2}^{2}+6\lambda_{2}\lambda_{3}+5\lambda_{3}^{2})&\dfrac{1}{2}(\lambda_{3}+2\lambda_{4})^{2}\\ 0&0&\dfrac{1}{2}(\lambda_{3}+2\lambda_{4})^{2}&\dfrac{3}{4}\lambda_{2}^{2}+(\lambda_{2}+\lambda_{3})^{2}\end{pmatrix}$ (221)

### v2

@@
 We compute the Annihilation Cross Section with Sommerfeld Enhancement and the Photon Flux arising from Dark Matter annihilation events.
-We found that the model exhibits a resonant mass part around the range of $5-10$ TeV.
+We found that the model exhibits a resonant mass part around the range of $2-10$ TeV.
 This resonance is influenced by the choice of portal Higgs coupling parameter.
-We show that these prediction can be partially tested by CTA in the near future
+We show that these prediction can be partially tested by CTA in the near future.
 1 Introduction
-Currently, one the most urgent problems in Particle Physics revolves around the elucidation of Dark Matter’s (DM) [2] fundamental nature.
+Currently, one of the most urgent problems in Particle Physics revolves around the elucidation of Dark Matter’s (DM) fundamental nature  [1].
 It is well known that, to account for the overarching phenomenology of DM, it suffices to posit the existence of a new particle falling within a mass range spanning from a few GeV to a few TeV, characterized by an interaction strength with the standard sector on the order of the weak interaction.
-This phenomenon is commonly referred to as the ”WIMP miracle,” with WIMP representing ”Weakly Interacting Massive Particle [12].
+This phenomenon is commonly referred to as the ”WIMP miracle”, with WIMP representing ”Weakly Interacting Massive Particle” [2].
 An intriguing realization of the WIMP idea is to postulate the DM as the neutral component of an electroweak multiplet.
-This paradigm, usually called “Minimal Dark Matter”, was proposed long ago [8] and have been studied in datail for scalar field [7] and fermionic fields [10].
+This paradigm, usually called “Minimal Dark Matter”, was proposed long ago [3] and have been studied in detail for scalar field [4] and fermionic fields [5].
 In the case of vector DM (VDM), an effective model including a VDM candidate originated from a vector field in the adjoint representation of $SU(2)_{L}$ was first proposed in [6] while an ultraviolet complete model was developed in [7].
-On the other had, in the fundamental representation, models including a vector DM candidate coming from an electroweak doublet were studied in  [15, 18].
+On the other hand, models including a vector DM candidate coming from an electroweak doublet were studied in  [8, 9].
 These vector models are consistent with experimental data, if the new fields has mass in the range of some few TeV.
@@
 In this circumstances, indirect search can be a privileged alternative to discover or rule out these models.
-However, it is well known that to obtain an accurate value of the DM annhihilation cross section in the non-relativistic regime, it is necessary to take into account non-perturbative effects like the so called Sommerfeld Enhancement (SE).
-The SE effect was computed for the cases of scalar DM field [3, 7, 9, 11, 13] showing that effect has a resonant part, fermion DM field [5, 10] in its doublet and triplet representation of $SU(2)_{L}$ and vector DM in the adjoint representation in  [1].
-In this work, we focus on computing the SE for the annihilation of VDM in the fundamental representation of $SU(2)_{L}$ and the photon flux (relevant for Indirect Detection  [6, 14, 19]) produced by pair annihilation of the VDM.
+However, it is well known that to obtain an accurate value of the DM annihilation cross section in the non-relativistic regime, it is necessary to take into account non-perturbative effects like the so called Sommerfeld Enhancement (SE).
+The SE effect was computed for the cases of scalar DM [4, 11, 12, 13, 14], fermion DM field [5, 15] in the doublet and triplet representation of $SU(2)_{L}$ and vector DM in the adjoint representation  [7].
+In this work, we focus on computing the SE for the annihilation of VDM in the fundamental representation of $SU(2)_{L}$ and the photon flux (relevant for Indirect Detection  [10, 16, 17, 18]) produced by pair annihilation of the VDM.
+As we will see in the following section, a unique feature of our case is the presence of non-minimal interactions between the massive vector doublet and the standard gauge bosons.
 Our work is organized as follows.
 In section 2, we describe the main aspects of the model.
-Section 3 is devoted to the computation of the annihilation matrices.
+Section 3 is devoted to the computation of the electroweak, annihilation matrices and annihilation cross section with SE.
 In section 4 we apply our results to compute the photon flux from DM annihilation at the center of the Milky Way and describe our results.
@@
 We start by recalling the main characteristics of the model initially developed in reference  [8].
-In this context, the Standard Model (SM) is extended by the introduction of a vector field $V_{\mu}$ which transforms as (1,2,1/2) under the standard group $SU(3)_{c}\times SU(2)_{L}\times U(1)_{Y}$ This new vector field can be explicitly written as:
+In this context, the Standard Model (SM) is extended by the introduction of a vector field $V_{\mu}$ which transforms as (1, 2, 1/2) under the standard group $SU(3)_{c}\times SU(2)_{L}\times U(1)_{Y}$.
+This new vector field can be explicitly written as:
 $V_{\mu}=\begin{pmatrix}V_{\mu}^{+}\\ V_{\mu}^{0}\end{pmatrix}=\begin{pmatrix}V_{\mu}^{+}\\ \dfrac{V_{\mu}^{1}+i~{}V_{\mu}^{2}}{\sqrt{2}}\end{pmatrix}$ (1)
 In principle, the two neutral states $V_{\mu}^{1}$ and $V_{\mu}^{2}$ (the vector and axial-vector fields) can have different masses.
-However, the requirement of unitarity preservation at low energies constrains the parameter space of the model in such a way that $V_{\mu}^{1}$ and $V_{\mu}^{2}$ become degenerated or nearly degenerated.
-In this work, we assume that $V_{\mu}^{1}$ and $V_{\mu}^{2}$ are exactly degenerated.
+However, unitarity constrains force the neutral components $V_{\mu}^{1}$ and $V_{\mu}^{2}$ to be fully degenerate  [8].
 In this case they combine into a complex neutral field $V_{\mu}^{0}$.
@@
 $+i\dfrac{g^{\prime}}{2}\kappa_{1}V_{\mu}^{\dagger}B^{\mu\nu}V_{\nu}+ig\kappa_{2}V_{\mu}^{\dagger}W^{\mu\nu}V_{\nu}$ (2)
-where $B^{\mu\nu}$ represents the Abelian strength field of $U(1)_{Y}$ and $W^{\mu\nu}=W^{\mu\nu~{}a}\dfrac{\tau^{a}}{2}$ represents the non-Abelian strength field of $SU(2)_{L}$.
-Notice that the las two terms of Lagrangian (2) describe non-minimal interactions between the new vector field and the gauge bosons.
+where $B^{\mu\nu}$ represents the Abelian strength field of $U(1)_{Y}$ and $W^{\mu\nu}=W^{\mu\nu a}\frac{\tau^{a}}{2}$ is the non-Abelian strength field of $SU(2)_{L}$.
+The lagrangian mass of the vector field $\tilde{m}$ will be corrected after the Higgs boson acquires its vacuum expectation value, due to the interaction terms between the Higgs field and the vector field present in the potential.
+We will call $m$ the physical mass of the vector boson.
+Notice that the last two terms of Lagrangian (2) describe non-minimal interactions between the new vector field and the gauge bosons.
+The appearance of this non-minimal dimension 4 terms is a special feature of having introduced a vector doublet.
+Similar non-minimal interactions in the scalar and fermion cases happen to be higher dimensional.
 In order to avoid a direct photon interaction with $V_{\mu}^{0}$, the otherwise arbitrary constants $\kappa_{1}$ and $\kappa_{2}$ have to be equal.
-In this work, however, for the sake of simplicity and minimality we assume that $\kappa_{1}=\kappa_{2}=0$.
-Besides, the coupling constant that regulates the interactions between the Higgs field and the new vector field can be treated a free parameters.
-For simplicity, we will take the assumption $\lambda_{2}=\lambda_{3}=\lambda_{4}=\lambda$).
+For the purpose of comparing, we work with two values of $\kappa=\kappa_{1}=\kappa_{2}$: $\kappa=0$ and $\kappa=1$.
+The last value coincides with the choice made in [8].
+Another important sector of the Lagrangian is the part of the potential which describes the interaction between the Higgs field and our massive vector doublet.
+Of course, the coupling constant that regulates the interactions between the Higgs field and the new vector field are free parameters.
+However, there is no reason to think that an important hierarchy arise among these couplings.
+Consequently, and for the sake of simplicity, we will take the assumption that all of them are of the same order $\lambda_{2}\approx\lambda_{3}\approx\lambda_{4}\approx\lambda$.
 In this way, $\lambda$ will be the only constant regulating the Higgs portal.
-Due to a fortunate conspiracy between the Lorentz and gauge structures of the model, the new vector field can only appear in pairs.
+Finally, the $V_{\mu}$ quartic self-interactions will not contribute significantly to the SE since this phenomenon is dominated by interactions mediated by particles much lighter than $V_{\mu}$, like the electroweak gauge bosons and the Higgs boson.
+Consequently, we can set $\alpha_{2}=\alpha_{3}=0$.
+Due to a fortunate conspiracy between the Lorentz and gauge structures of the model, the new vector field can only appear in pairs in the Lagrangian.
 That means that the Lagrangian is invariant by a $Z_{2}$ transformation under which the new vector is odd and all the standard fields are even.
@@
 This effect may produce sensible consequences for the annihilation cross section in the non-relativistic regime [5, 19].
-In the few next paragraphs we will review the formalism used to compute the Sommerfeld enhancement effect and the we will apply it to our model.
+In the next few paragraphs we will review the formalism used to compute the Sommerfeld enhancement effect and the we will apply it to our model.
 We start by considering the DM annihilation processes, which are described by a two-body state, in the non-relativistic regime.
 The non-relativistic Schwinger-Dyson equation for the Green function $\bra{0}T\phi_{b}(R^{\prime},\textbf{r}^{\prime})\phi_{a}^{\dagger}(R,\textbf{r})\ket{0}$ can be written as:
-$\left[\left(i\partial_{R^{0}}+\frac{\nabla_{R}^{2}}{4m}+\frac{\nabla_{r}^{2}}{m}\right)-V(r)+i\frac{N}{2}\Gamma\delta^{3}(r)\right]$ $\bra{0}T\phi_{b}(R^{\prime},\textbf{r}^{\prime})\phi_{a}^{\dagger}(R,\textbf{r})\ket{0}$ (7)
+$\left[\left(i\partial_{R^{0}}+\frac{\nabla_{R}^{2}}{4M_{V}}+\frac{\nabla_{r}^{2}}{M_{V}}\right)-V(r)+i\frac{N}{2}\Gamma\delta^{3}(r)\right]$ $\bra{0}T\phi_{b}(R^{\prime},\textbf{r}^{\prime})\phi_{a}^{\dagger}(R,\textbf{r})\ket{0}$ (3)
 $=i\delta^{4}(R-R^{\prime})\delta^{3}(\textbf{r}-\textbf{r}^{\prime})\delta_{ab}$
-Here, $m$ represents the mass of Dark Matter, $\phi_{i}$ represents the DM field, $N$ is the number spin states for the two bodies system (for spin 1, we have $N=9$ ) and $\Gamma$ is the so called annihilation matrix.
+Here, $m$ represents the mass of Dark Matter, $\phi_{i}$ represents the DM field, $N$ is the number spin states for the two bodies system (for spin 1, we have $N=9$), $V(r)$ is electroweak matrices and $\Gamma$ is the so called annihilation matrix.
 The coordinate $=(R^{0},\textbf{R})$ R ) denotes the center-of-mass coordinate, and r represents the relative coordinate of the Dark Matter multiplets, defined as:
@@
 The imaginary part of the diagonal component of this Green function $\bra{0}T\phi_{b}\phi_{a}^{\dagger}\ket{0}$, is associated with the total annihilation cross section.
-We proceed to decompose the Green functions into individual partial wave modes
+We proceed to decompose the Green functions into individual partial wave modes.
 $\bra{0}T\phi_{b}(R^{\prime},\textbf{r}^{\prime})\phi_{a}^{\dagger}(R,\textbf{r})\ket{0}=\int\dfrac{d^{4}P}{(2\pi)^{4}}e^{-iP\cdot(R-R^{\prime})}\sum_{l}\dfrac{2l+1}{4\pi}P_{l}(\cos\gamma)(-i)G_{ba}^{(E,l)}(r,r^{\prime})$ (5)
 where $\gamma$ denotes the relative angle between r and $\textbf{r}^{\prime}$. The non-relativistic energy for a dark matter pair ($E$) is
-$E=P^{0}-\dfrac{\textbf{P}}{4m}=\dfrac{mv^{2}}{4}$ (10)
+$E=P^{0}-\dfrac{\textbf{P}^{2}}{4M_{V}}=\dfrac{M_{V}v^{2}}{4}$ (6)
 where, $v$ represents the relative velocity between a pair of dark matter particles. By substituting (3) into (5), we obtain the following result:
-$\left[-\dfrac{1}{mr}\dfrac{d^{2}}{dr^{2}}r-\dfrac{l(l+1)}{mr^{2}}-E+V(r)-i\Gamma\dfrac{N\delta(r)}{8\pi r^{2}}\right]G_{ba}^{(E,l)}(r,r^{\prime})=\dfrac{\delta(r-r^{\prime})}{r^{2}}\delta_{ba}$ (11)
+$\left[-\dfrac{1}{M_{V}r}\dfrac{d^{2}}{dr^{2}}r-\dfrac{l(l+1)}{M_{V}r^{2}}-E+V(r)-i\Gamma\dfrac{N\delta(r)}{8\pi r^{2}}\right]G_{ba}^{(E,l)}(r,r^{\prime})=\dfrac{\delta(r-r^{\prime})}{r^{2}}\delta_{ba}$ (7)
 In this analysis, we focus on the annihilation processes with orbital angular momentum $l=0$ (s-wave), which provide the leading-order contributions in the non-relativistic limit.
-Consequently, the total angular momentum $J$ is equal to the total spin angular momentum $S$.
+Indeed, we verified that the $l=1$ contribution is at least one order of magnitude smaller than the leading s-wave.
+Consequently, in our calculations the total angular momentum $J$ is equal to the total spin angular momentum $S$.
 The leading-order potential terms exhibit spherical symmetry, resulting in the conservation of $J$ in annihilation processes.
@@
 To simplify the differential equation, we introduce $G_{ba}^{(E,l)}(r,r^{\prime})=\frac{g_{ba}(r,r^{\prime})}{rr^{\prime}}$.
-$\left[-\dfrac{1}{mr}\dfrac{d^{2}}{dr^{2}}r-E+V(r)-i\Gamma\dfrac{N\delta(r)}{8\pi r^{2}}\right]g_{ba}^{(E,l)}(r,r^{\prime})=\delta(r-r^{\prime})\delta_{ba}$ (12)
-The cross-section is determined by the imaginary part of the solution, and we derive the leading-order solution within the perturbative expansion of $\alpha$.
+$\left[-\dfrac{1}{M_{V}r}\dfrac{d^{2}}{dr^{2}}r-E+V(r)-i\Gamma\dfrac{N\delta(r)}{8\pi r^{2}}\right]g_{ba}^{(E,l=0)}(r,r^{\prime})=\delta(r-r^{\prime})\delta_{ba}$ (8)
+The cross-section is determined by the imaginary part of the solution, and we derive the leading-order solution within the perturbative expansion of $\alpha_{2}$.
 The leading order solution can be written as
@@
 In the above expression, we introduce $g_{>ab}$ which is the solution of the following equation
-$\left[-\dfrac{1}{m}\dfrac{d}{dr^{2}}-E+V(r)\right]g_{>ab}(r)=0$ (14)
-with the boundary conditions
+$\left[-\dfrac{1}{M_{V}}\dfrac{d}{dr^{2}}-E+V(r)\right]g_{>ab}(r)=0$ (10)
+Where $V(r)=V_{g-V}(r)+V_{h-V}(r)$ represents the electroweak potential, composed of $V_{g-V}(r)$ (interaction gauge bosons-VDM) and $V_{h-V}(r)$ (interaction Higgs boson-VDM) terms, $E$ denotes the energy of dark matter particles, $m$ stands for the mass of dark matter, and $g_{>ab}(r)$ denotes the wave function for $(r\gg 1)$, with the corresponding boundary conditions.
 $g_{>ab}(r=0)=\delta_{ba}$ (11)
@@
 In particular, we establish the following condition to govern the asymptotic behavior.
-$g_{>ba}(r)\longrightarrow\ \left\{\begin{array}[]{lcc}d_{ba}(E_{a})e^{i\sqrt{mE_{b}}r}\hskip 28.45274ptshort-range~{}potential\\ \\ d_{ba}(E_{a})e^{i\sqrt{mE_{b}}r}e^{i\frac{m\alpha}{2\sqrt{mE_{b}}}\log{2\sqrt{mE_{b}}r}}\hskip 28.45274ptlarge-range~{}potential\\ \end{array}\right.$ (19)
+$g_{>ba}(r)\longrightarrow\ \left\{\begin{array}[]{lcc}d_{ba}(E_{a})e^{i\sqrt{M_{V}E_{b}}r}\hskip 28.45274ptshort-range~{}potential\\ \\ d_{ba}(E_{a})e^{i\sqrt{M_{V}E_{b}}r}e^{i\frac{M_{V}\alpha}{2\sqrt{M_{V}E_{b}}}\log{2\sqrt{M_{V}E_{b}}r}}\hskip 28.45274ptlarge-range~{}potential\\ \end{array}\right.$ (15)
 Here, $E_{a}$ and $E_{b}$ denote the non-relativistic kinetic energy of $\phi_{a}$ and $\phi_{b}$, respectively.
@@
 We numerically solve equation (10) to obtain the Sommerfeld factor as a function of the relative velocity for the initial state $v$.
-This effect becomes significant in the non-relativistic regime ($v\ll 1$), where the potential energy becomes comparable to the kinetic energy ($E\simeq\frac{mv^{2}}{4}$ ) in the Schrödinger equations.
-Conversely, the enhancement factors approach unity for $v\gg 1$.
+This effect becomes significant in the non-relativistic regime ($v\ll 1$), where the potential energy becomes comparable to the kinetic energy ($E\simeq\frac{M_{V}v^{2}}{4}$ ) in the Schrödinger equations.
+Conversely, the enhancement factors approach unity for ($v\approx 1$).
 Consequently, we can recover the perturbative cross-section, as obtained by the Born-Oppenheimer approximation, by replacing the Sommerfeld factor with $1$.
@@
 If we substitute the leading order solution given in (9), we obtain
-$\braket{\sigma_{a}v}=c_{a}\dfrac{m^{2}}{k^{2}}\lim_{E\longrightarrow\frac{k^{2}}{m^{2}}}\left(E-\dfrac{k^{2}}{m^{2}}\right)^{2}[\mathcal{A}(k)~{}\Gamma~{}\mathcal{A}^{\dagger}(k)]_{aa}$ (25)
+$\braket{\sigma_{a}v}=c_{a}\dfrac{M_{V}^{2}}{k^{2}}\lim_{E\longrightarrow\frac{k^{2}}{M_{V}^{2}}}\left(E-\dfrac{k^{2}}{m^{2}}\right)^{2}[\mathcal{A}(k)~{}\Gamma~{}\mathcal{A}^{\dagger}(k)]_{aa}$ (21)
 where we define
@@
 […]
-$=d_{ba}(E)\left(-\dfrac{k}{m}\right)\left(E-\frac{k^{2}}{m}\right)^{-1}$ (28)
+$=d_{ba}(E)\left(-\dfrac{k}{M_{V}}\right)\left(E-\frac{k^{2}}{M_{V}}\right)^{-1}$ (24)
 We replace into (21), we express the annihilation cross section as below
@@
 $\braket{\sigma_{a}v}_{XX^{\prime}}=c\sum_{\alpha,\beta}\sum_{J,J_{z}}(\Gamma_{XX^{\prime}}^{J,J_{z}})_{\alpha\beta}~{}d_{3\alpha}(E)d_{3\beta}^{\ast}(E)$ (26)
-Here, $c$ is set to $2$ for initial states consisting of identical particles.
-We define $E$ as approximately $E\simeq\frac{mv_{rel}^{2}}{4}$ representing the non-relativistic kinetic energy of the V-particles, and $d_{\alpha\beta}(E)~{}(\alpha,\beta=1,2,3,4)$ as the Sommerfeld enhancement factor.
+Here, the index $3$ must be fixed; therefore, the Sommerfeld enhancement factor implies that we are considering the process $(V_{0}V_{0}^{\ast}\longrightarrow XX^{\prime})$, with $c$ is set to $2$ for initial states consisting of identical particles.
+We define $E$ as approximately $(E\simeq\frac{M_{V}v_{rel}^{2}}{4})$ representing the non-relativistic kinetic energy of the V-particles, $(\Gamma_{XX^{\prime}}^{J,J_{z}})_{\alpha\beta}$ is the Annihilation matrices, with $X$ is any standard model particle, and $d_{\alpha\beta}(E)$ with $(\alpha,\beta=1,2,3,4)$ as the Sommerfeld enhancement factor.
 3.3 Non-Relativistic Limit
 To ensure that the dark matter vector remains non-relativistic, we must carry out the following non-relativistic limit of the operators, as defined in [7].
-$V_{\mu}^{0}(x)=\dfrac{1}{\sqrt{2m}}[v_{0}(\vec{x})\epsilon_{\mu}(x)e^{imt}+a_{0}^{\dagger}(\vec{x})\epsilon_{\mu}^{\ast}(x)e^{-imt}]\hskip 14.22636pt$ (31)
-$(V_{\mu}^{0})^{\ast}(x)=\dfrac{1}{\sqrt{2m}}[a_{0}(\vec{x})\epsilon_{\mu}(x)e^{imt}+v_{0}^{\dagger}(\vec{x})\epsilon_{\mu}^{\ast}(x)e^{-imt}]\hskip 14.22636pt$ (32)
-$V_{\mu}^{+}(x)=\dfrac{1}{\sqrt{2m}}[v_{+}(\vec{x})\epsilon_{\mu}(x)e^{imt}+v_{-}^{\dagger}(\vec{x})\epsilon_{\mu}^{\ast}(x)e^{-imt}]\hskip 14.22636pt$ (33)
-[…]
+$V_{\mu}^{0}(x)=\dfrac{1}{\sqrt{2M_{V}}}[v_{0}(\vec{x})\epsilon_{\mu}(x)e^{iM_{V}t}+a_{0}^{\dagger}(\vec{x})\epsilon_{\mu}^{\ast}(x)e^{-iM_{V}t}]\hskip 14.22636pt$ (27)
+$(V_{\mu}^{0})^{\ast}(x)=\dfrac{1}{\sqrt{2M_{V}}}[a_{0}(\vec{x})\epsilon_{\mu}(x)e^{iM_{V}t}+v_{0}^{\dagger}(\vec{x})\epsilon_{\mu}^{\ast}(x)e^{-iM_{V}t}]\hskip 14.22636pt$ (28)
+$V_{\mu}^{+}(x)=\dfrac{1}{\sqrt{2M_{V}}}[v_{+}(\vec{x})\epsilon_{\mu}(x)e^{iM_{V}t}+v_{-}^{\dagger}(\vec{x})\epsilon_{\mu}^{\ast}(x)e^{-iM_{V}t}]\hskip 14.22636pt$ (29)
+$V_{\mu}^{-}(x)=\dfrac{1}{\sqrt{2M_{V}}}[v_{+}^{\dagger}(\vec{x})\epsilon_{\mu}^{\ast}(x)e^{-iM_{V}t}+v_{-}(\vec{x})\epsilon_{\mu}(x)e^{iM_{V}t}]\hskip 14.22636pt$ (30)
 Here, $v_{0},a_{0},v_{+},v_{-}$ represent the non-relativistic limits of the dark matter vector field.
@@
 Specifically, we consider the pairs ($v_{0}v_{0}$), for which such annihilation occurs.
-The most important contributions come from processes involving the exchange of $Z$ and $W^{\pm}$ bosons, having in their initial states ($v_{0}v_{0}$) , ($v_{0}a_{0}$), or ($v_{+}v_{-}$) pairs.
+The most important contributions come from processes involving the exchange of $\gamma$, $Z$, $W^{\pm}$ bosons, having in their initial states ($v_{0}v_{0}$) , ($v_{0}a_{0}$), or ($v_{+}v_{-}$) pairs.
 Our approach involves deriving an effective action that accurately describes these pairs.
@@
 From these expressions, we employ the non-relativistic limit to obtain the electroweak matrix, which we will use to solve the Schrödinger equation (10).
-In the non-relativistic limit, the first term of equation (3.4.1) characterizes the Coulomb interaction between non-relativistic particles, whereas the remaining terms depict the Yukawa potentials related to the exchange of $Z$ and $W$ bosons.
+In the non-relativistic limit, the first term of equation (3.4.1) characterizes the Coulomb potential between non-relativistic particles, whereas the remaining terms depict the Yukawa potentials related to the exchange of $Z$ and $W$ bosons.
 Upon employing vector (32), the effective action transforms into the following expression:
-$S_{eff,1_{NR}}=-\int d^{4}xd^{3}y~{}s(x,\vec{y})^{\dagger}V_{Gauge}(|\vec{x}-\vec{y}|)s(x,\vec{y})$ (42)
-where $V_{Gauge}(|\vec{x}-\vec{y}|)$ is electroweak potential matrix (interaction gauge boson-vector dark matter), with $r=|\vec{x}-\vec{y}|$;
-$V_{Gauge}(r)=-\dfrac{g^{2}}{8\pi r}\begin{pmatrix}4s_{W}^{2}e^{-M_{Z}r}&0&0&0\\ 0&4s_{W}^{2}e^{-M_{Z}r}&0&0\\ 0&0&16s_{W}^{2}e^{-M_{Z}r}&2\sqrt{2}e^{-M_{W}r}\\ 0&0&2\sqrt{2}e^{-M_{W}r}&8s_{W}^{2}e^{-M_{Z}r}-2(c_{W}-s_{W}t_{W}^{2})^{2}\end{pmatrix}$ (47)
-In appendix A we show how to compute this matrix.
+$S_{eff,1_{NR}}=-\int d^{4}xd^{3}y~{}s(x,\vec{y})^{\dagger}V_{g-V}(|\vec{x}-\vec{y}|)s(x,\vec{y})$ (34)
+where $V_{g-V}(|\vec{x}-\vec{y}|)$ is electroweak potential matrix (interaction gauge boson-vector dark matter), with $r=|\vec{x}-\vec{y}|$;
+$V_{g-V}(r)=-\dfrac{g^{2}}{8\pi r}\begin{pmatrix}4a^{2}e^{-M_{Z}r}&0&0&0\\ 0&4a^{2}e^{-M_{Z}r}&0&0\\ 0&0&16a^{2}e^{-M_{Z}r}&b^{2}e^{-M_{W}r}\\ 0&0&b^{2}e^{-M_{W}r}&8a^{2}e^{-M_{Z}r}-2(c_{W}(1-(\kappa_{1}+\kappa{2})-s_{W}t_{W}^{2})^{2}\end{pmatrix}$ (35)
+with $a=(c_{W}(\kappa_{1}+\kappa_{2})+s_{W})$, $b=(\frac{3}{8}-2\sqrt{2})(\kappa_{1}+\kappa_{2})+2\sqrt{2}$ In appendix A we show how to compute this matrix.
 3.4.2 Higgs Boson-Dark Matter Interaction
@@
 Our primary focus lies in understanding the cubic interaction, as it is crucial for unraveling the nature of this interaction.
-For this purpose, we parametrize the Higgs doublet as folloes:
+For this purpose, we parameterize the Higgs doublet as follows:
 $\phi(x)=\begin{pmatrix}\phi^{+}(x)\\ \dfrac{v+h(x)}{\sqrt{2}}+i\phi^{0}(x)\end{pmatrix}$ (36)
-where $\phi^{+}$ and $\phi^{0}$ are would-be Goldtone bosons, and $v$ is vaccum expectation value.
+where $\phi^{+}$ and $\phi^{0}$ are would-be Goldtone bosons, and $v$ is vacuum expectation value.
 Now, when we substitute $V_{\mu}$, $V_{\mu}^{\dagger}$, $\phi$, and $\phi^{\dagger}$ into the Lagrangian (2), we can derive the electroweak matrix describing the interaction between the Higgs and the Dark Matter vector, which takes the form:
-$V_{h-V}(r)=\dfrac{v^{2}}{32\pi m^{2}r}\begin{pmatrix}8\lambda_{4}^{2}a&0&0&0\\ 0&8\lambda_{4}^{2}a&0&0\\ 0&0&\dfrac{1}{2}(\lambda_{2}+\lambda_{3})^{2}a&(\lambda_{3}+2\lambda_{4})^{2}e^{-M_{W}r}\\ 0&0&(\lambda_{3}+2\lambda_{4})^{2}e^{-M_{W}r}&\lambda_{2}^{2}a\end{pmatrix}$ (55)
-In Appendix A, we illustrate the methodology for computing this matrix.
+$V_{h-V}(r)=\dfrac{v^{2}}{32\pi M_{V}^{2}r}\begin{pmatrix}8\lambda_{4}^{2}a&0&0&0\\ 0&8\lambda_{4}^{2}a&0&0\\ 0&0&\dfrac{1}{2}(\lambda_{2}+\lambda_{3})^{2}a&(\lambda_{3}+2\lambda_{4})^{2}e^{-M_{W}r}\\ 0&0&(\lambda_{3}+2\lambda_{4})^{2}e^{-M_{W}r}&\lambda_{2}^{2}a\end{pmatrix}$ (37)
+with $a=(2e^{-M_{h}r}+e^{-M_{Z}r})$. In Appendix A, we illustrate the methodology for computing this matrix.
 3.5 Annihilation Matrix
@@
 In Appendix B, the computation of these terms is detailed. Considering the different channels and total spin states, we obtain the following annihilation matrices:
-$\Gamma_{WW}^{J=0}=\dfrac{2g^{4}}{64\sqrt{3}\pi m^{2}}\begin{pmatrix}2&0&0&0\\ 0&2&0&0\\ 0&0&1&1\\ 0&0&1&1\end{pmatrix}$ (61)
-$\Gamma_{WW}^{J=2}=\dfrac{g^{4}}{64\pi m^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2&0&0&0\\ 0&2&0&0\\ 0&0&1&1\\ 0&0&1&1\end{pmatrix}$ (66)
-$\Gamma_{AA}^{J=0}=\dfrac{e^{4}}{128\sqrt{3}\pi m^{2}}\begin{pmatrix}2a^{2}&0&0&0\\ 0&2a^{2}&0&0\\ 0&0&a^{2}&4a^{2}\\ 0&0&4a^{2}&4\end{pmatrix}$ (71)
-$\Gamma_{AA}^{J=2}=\dfrac{e^{4}}{256\pi m^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2a^{2}&0&0&0\\ 0&2a^{2}&0&0\\ 0&0&a^{2}&4a^{2}\\ 0&0&4a^{2}&4\end{pmatrix}$ (76)
+$\Gamma_{WW}^{J=0}=\dfrac{2g^{4}}{64\sqrt{3}\pi M_{V}^{2}}\begin{pmatrix}2&0&0&0\\ 0&2&0&0\\ 0&0&1&1\\ 0&0&1&1\end{pmatrix}$ (39)
+$\Gamma_{WW}^{J=2}=\dfrac{g^{4}}{64\pi M_{V}^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2&0&0&0\\ 0&2&0&0\\ 0&0&1&1\\ 0&0&1&1\end{pmatrix}$ (40)
+$\Gamma_{AA}^{J=0}=\dfrac{e^{4}}{128\sqrt{3}\pi M_{V}^{2}}\begin{pmatrix}2a^{2}&0&0&0\\ 0&2a^{2}&0&0\\ 0&0&a^{2}&4a^{2}\\ 0&0&4a^{2}&4\end{pmatrix}$ (41)
+$\Gamma_{AA}^{J=2}=\dfrac{e^{4}}{256\pi M_{V}^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2a^{2}&0&0&0\\ 0&2a^{2}&0&0\\ 0&0&a^{2}&4a^{2}\\ 0&0&4a^{2}&4\end{pmatrix}$ (42)
 with $a=(1-\tan\theta_{W})$
-$\Gamma_{ZZ}^{J=0}$ $=\dfrac{g^{4}}{128\sqrt{3}\pi m^{2}}\begin{pmatrix}2b^{2}&0&0&0\\ 0&2b^{2}&0&0\\ 0&0&2b^{2}&-4c\\ 0&0&-4c&4\sin^{4}\theta_{W}\end{pmatrix}$ (81)
-$\Gamma_{ZZ}^{J=2}$ $=\dfrac{g^{4}}{256\pi m^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2b^{2}&0&0&0\\ 0&2b^{2}&0&0\\ 0&0&2b^{2}&-4c\\ 0&0&-4c&4\sin^{4}\theta_{W}\end{pmatrix}$ (86)
+$\Gamma_{ZZ}^{J=0}$ $=\dfrac{g^{4}}{128\sqrt{3}\pi M_{V}^{2}}\begin{pmatrix}2b^{2}&0&0&0\\ 0&2b^{2}&0&0\\ 0&0&2b^{2}&-4c\\ 0&0&-4c&4\sin^{4}\theta_{W}\end{pmatrix}$ (43)
+$\Gamma_{ZZ}^{J=2}$ $=\dfrac{g^{4}}{256\pi M_{V}^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2b^{2}&0&0&0\\ 0&2b^{2}&0&0\\ 0&0&2b^{2}&-4c\\ 0&0&-4c&4\sin^{4}\theta_{W}\end{pmatrix}$
 with $b=(\tan^{2}\theta_{W}\sin^{2}\theta_{W}-\sin\theta_{W}\cos\theta_{W})$ y $c=\sin^{2}\theta_{W}\tan\theta_{W}(\sin^{2}\theta_{W}-\cos^{2}\theta_{W})$
-$\Gamma_{AZ}^{J=0}$ $=\dfrac{e^{2}g^{2}}{128\sqrt{3}\pi m^{2}}\begin{pmatrix}d^{2}&0&0&0\\ 0&d^{2}&0&0\\ 0&0&d^{2}&f\\ 0&0&f&\dfrac{\sin^{2}(2\theta_{W})}{\cos^{2}\theta_{W}}\end{pmatrix}$ (91)
-$\Gamma_{AZ}^{J=2}$ $=\dfrac{e^{2}g^{2}}{256\pi m^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}d^{2}&0&0&0\\ 0&d^{2}&0&0\\ 0&0&d^{2}&f\\ 0&0&f&\dfrac{\sin^{2}(2\theta_{W})}{\cos^{2}\theta_{W}}\end{pmatrix}$ (96)
+$\Gamma_{AZ}^{J=0}$ $=\dfrac{e^{2}g^{2}}{128\sqrt{3}\pi M_{V}^{2}}\begin{pmatrix}d^{2}&0&0&0\\ 0&d^{2}&0&0\\ 0&0&d^{2}&f\\ 0&0&f&\dfrac{\sin^{2}(2\theta_{W})}{\cos^{2}\theta_{W}}\end{pmatrix}$ (44)
+$\Gamma_{AZ}^{J=2}$ $=\dfrac{e^{2}g^{2}}{256\pi M_{V}^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}d^{2}&0&0&0\\ 0&d^{2}&0&0\\ 0&0&d^{2}&f\\ 0&0&f&\dfrac{\sin^{2}(2\theta_{W})}{\cos^{2}\theta_{W}}\end{pmatrix}$
 with $d=(\cos\theta_{W}-\sin\theta_{W})$ y $f=(\cos\theta_{W}-\sin\theta_{W})^{4}(1+\tan\theta_{W})^{2}$
@@
 The resulting annihilation matrix can be written as follows:
-$\Gamma_{Higgs}=\dfrac{1}{32\pi m^{2}}\begin{pmatrix}\left(\dfrac{3}{2}+\sqrt{2}\right)^{2}\lambda_{4}^{2}&0&0&0\\ 0&\dfrac{9}{4}\lambda_{4}^{2}&0&0\\ 0&0&\dfrac{1}{4}(7\lambda_{2}^{2}+6\lambda_{2}\lambda_{3}+5\lambda_{3}^{2})&\dfrac{1}{2}(\lambda_{3}+2\lambda_{4})^{2}\\ 0&0&\dfrac{1}{2}(\lambda_{3}+2\lambda_{4})^{2}&\dfrac{3}{4}\lambda_{2}^{2}+(\lambda_{2}+\lambda_{3})^{2}\end{pmatrix}$ (101)
-3.6 Results
+$\Gamma_{Higgs}=\dfrac{1}{32\pi M_{V}^{2}}\begin{pmatrix}\left(\dfrac{3}{2}+\sqrt{2}\right)^{2}\lambda_{4}^{2}&0&0&0\\ 0&\dfrac{9}{4}\lambda_{4}^{2}&0&0\\ 0&0&\dfrac{1}{4}(7\lambda_{2}^{2}+6\lambda_{2}\lambda_{3}+5\lambda_{3}^{2})&\dfrac{1}{2}(\lambda_{3}+2\lambda_{4})^{2}\\ 0&0&\dfrac{1}{2}(\lambda_{3}+2\lambda_{4})^{2}&\dfrac{3}{4}\lambda_{2}^{2}+(\lambda_{2}+\lambda_{3})^{2}\end{pmatrix}$ (45)
+3.6 Numerical Results
 Now, we have all the needed ingredients to compute the thermal averaged cross-section $\braket{\sigma v}$ using equation (26).
@@
 $\braket{\sigma v}=\braket{\sigma v}_{W^{+}W^{-}}+\braket{\sigma v}_{ZZ}+\braket{\sigma v}_{\gamma\gamma}+\braket{\sigma v}_{Z\gamma}+\braket{\sigma v}_{hh}.$ (46)
-Our results are shown in figure 2 .
-There are several noteworthy characteristics of the Sommerfeld enhancement that merit discussion.
-Firstly, it is a velocity-dependent phenomenon.
-For instance, no structure is observed for a relative velocity of $v=0.3~{}c$ (dashed lines in figure 2), which is representative of the early Universe, when the relic density was formed.
-In contrast, for smaller velocities, like the one expected in the galaxy ($v=10^{-3}c$), significant oscillations and resonances appear.
-For masses that exhibit such resonances (roughly around $6$ TeV), the Sommerfeld factor experiences a significant boost of about 5 order of magnitude, for all the values of portal coupling constant $\lambda$.
-Figure 2: Annihilation cross-section (fundamental representation spin-1 of $SU(2)_{L}$) for the process $v_{0}a_{0}\longrightarrow SMSM$ with the Sommerfeld Enhancement factor, considering velocities $v=10^{-3}$ (galactic Halo, shown as continuous lines, green, blue, black and gray lines ) and $v=0.3$ (Relic density, indicated by dashed lines), and different values of the Higgs portal coupling constant $\lambda$.
+Our results are shown in figures 2 and 3 .
+The first one shows our predictions without taking into account the non-minimal interactions ($\kappa=0$) while in the second figure we used $\kappa=1$.
+In both cases, there are several noteworthy characteristics of the Sommerfeld enhancement that merit discussion.
+Firstly, as usual, the SE is a velocity-dependent phenomenon.
+For instance, no structure is observed for a relative velocity of $v=0.3$ (dashed lines in figures 2 and 3), which is representative of the early Universe, when the relic density was formed.
+In contrast, for smaller velocities, like the one expected in the galaxy ($v=10^{-3}$), significant oscillations and resonances appear.
+We have observed that the small oscillations dependent on interactions mediated by the Higgs boson.
+Indeed, we see that they are attenuated for large values of $\lambda$ parameter Interestingly, the presence of the non-minimal interactions introduce an additional resonance at a lower mass.
+However, the $\kappa$-terms don’t modify significantly the size of the oscillations nor the amplitude of the resonances.
+For masses that exhibit such resonances (roughly around $1.9$ and $6$ TeV), the Sommerfeld factor experiences a significant boost of about 5 order of magnitude, for all the values of portal coupling constant $\lambda$.
+Figure 2: Annihilation cross-section (fundamental representation spin-1 of $SU(2)_{L}$) for the process $v_{0}a_{0}\longrightarrow SMSM$ with the Sommerfeld Enhancement factor, considering velocities $v=10^{-3}$ (galactic Halo, shown as continuous lines, green, blue, black and gray lines ) and $v=0.3$ (Relic density, indicated by dashed lines), and different values of the Higgs portal coupling constant $\lambda$, with $\kappa_{1}=\kappa_{2}=0$.
+The green line indicates that it is the spin-1 triplet representation of $SU(2)_{L}$, while the red line indicates that it is the spin-1/2 doublet representation of $SU(2)_{L}$.
+Figure 3: Annihilation cross-section (fundamental representation spin-1 of $SU(2)_{L}$) for the process $v_{0}a_{0}\longrightarrow SMSM$ with the Sommerfeld Enhancement factor, considering velocities $v=10^{-3}$ (galactic Halo, shown as continuous lines, green, blue, black and gray lines ) and $v=0.3$ (Relic density, indicated by dashed lines), and different values of the Higgs portal coupling constant $\lambda$, with $\kappa_{1}=\kappa_{2}=1$.
 The green line indicates that it is the spin-1 triplet representation of $SU(2)_{L}$, while the red line indicates that it is the spin-1/2 doublet representation of $SU(2)_{L}$.
@@
 The $\gamma$-ray flux resulting from the annihilation of neutral Dark Matter particles, as described in [17], in a region of the sky with a solid angle $\Omega$ is given by:
-$\dfrac{d\Phi}{dE}=\dfrac{1}{8\pi}J(\Omega)\dfrac{\braket{\sigma v}_{\gamma\gamma}}{m^{2}}\left(\dfrac{dN}{dE}\right)_{\gamma\gamma}$ (103)
+$\dfrac{d\Phi}{dE_{\gamma}}=\sum_{i}\dfrac{1}{8\pi}J(\Omega)\dfrac{\braket{\sigma v}_{i}}{M_{V}^{2}}\left(\dfrac{dN_{i}}{dE_{\gamma}}\right)$ (47)
 Here, $J(\Omega)$ represents the astrophysical $J$-factor, which measures the relevance and interest of the astrophysical target for Dark Matter annihilation.
 […]
-The $J$-factors for dwarf galaxies are approximately $J\approx 10^{19}$ $GeV^{2}/cm^{5}$, for the Andromeda galaxy, $J\approx 10^{20}$ $GeV^{2}/cm^{5}$, and for the Milky Way’s Galactic Center, $J\approx 10^{25}$ $GeV^{2}/cm^{5}$.
-If the solid angle changes by one degree, the $J$-factor undergoes an approximate change of $J\approx 10^{24}$ $GeV^{2}/cm^{5}$.
-The distribution, denoted as $\left(\frac{dN}{dE}\right)_{\gamma\gamma}$, characterizes the number of photons with energy $E_{\gamma\gamma}=m$, where $m$ represents the mass of the WIMP.
-The distribution is defined as follows:
-$\left(\dfrac{dN}{dE}\right)_{\gamma\gamma}=2\times\delta(E-E_{\gamma\gamma})$ (105)
-while $\braket{\sigma v}_{\gamma\gamma}$ represents the annihilation cross-section rate with the mediator, which is dependent on the branching ratio, and is defined as $\braket{\sigma v}_{\gamma\gamma}=BR_{\gamma\gamma}\times\braket{\sigma v}_{XX^{\prime}}$.
+where $\rho[r(\ell,\psi)]$ is the DM density and depends of the assumed astrophysical profile.
+The $J$-factor for dwarf galaxies is approximately $J\approx 10^{19}$ GeV2/cm5 while for the Andromeda galaxy, $J\approx 10^{20}$ GeV2/cm5, and for the Milky Way’s Galactic Center, $J\approx 10^{25}$ GeV2/cm5.
+If the solid angle changes by one degree, the $J$-factor undergoes an approximate change of $J\approx 10^{24}$ GeV2/cm5 [20].
+The distribution, denoted as $\left(\frac{dN_{i}}{dE_{\gamma}}\right)$, characterizes the number of photons with energy $E_{\gamma}$, while $\braket{\sigma v}_{i}$ represents the averaged annihilation cross-section from the i-th channel of annihilation.
+This quantity dependents on the branching ratio, and is defined as $\braket{\sigma v}_{\gamma\gamma}=BR_{\gamma\gamma}\times\braket{\sigma v}_{XX^{\prime}}$.
 Here, $\braket{\sigma v}_{XX^{\prime}}$ corresponds to the total cross-section for Dark Matter annihilation, as shown in equation (26).
-$BR_{\gamma\gamma}=\dfrac{\Gamma_{\gamma\gamma}}{\sum_{V}\Gamma_{VV^{\prime}}}$ (106)
-where $V$ is the gauge and Higgs bosons. Branching ratio is the quotient between annihilation matrix of photons by add total annihilation matrix.
-4.1 Results
-Using the expressions above, we calculate the photon flux resulting from the Annihilation Cross Section (30), which serves as a method for the Indirect Detection of Dark Matter.
-This entails the integration of equation (103) using the distribution of the number of photons as defined in (105).
-Figure 3 depict the photon flux from VDM annihilation per year.
-We can see, as expected, that as the Higgs portal increases in size, the photon flux also increases.Our constructions are not constrained by Fermi-LAT, as our region of interest is above $100$ GeV which is not tested by Fermi-LAT; therefore, in the graphs, we compare with the projected CTA sensitivity [16].
-We can see that for low vlues of the Higgs-DM coupling CTA can test only the region around the big resonance while for larger values CTA will be able to observe or rule out all our region of interest.
-Figure 3: Photons Flux from Annihilation of neutral Dark Matter particles (Yellow, Dark Orange, Purple and Dark Pink lines correspond to fundamental representation of $SU(2)_{L}$ spin-1), black line is the projected CTA photon flux sensitivity, The green line corresponds to the photon flux of the spin-1 $SU(2)_{L}$ triplet representation, while the red line corresponds to the $SU(2)_{L}$ doublet representation for spin-1/2.
+Using the expressions above, we calculate the photon flux resulting from the annihilation cross section (26), which serves as a method for the indirect detection of DM.
+This entails the integration of equation (47).
+Figures 4 and 5 depict the photon flux from VDM annihilation per second for $\kappa=0$ and $\kappa=1$ respectively.
+In all the cases we have used $\lambda=0.5$.
+We present the signals of photon flux resulting from the annihilation of VDM particles from the center of the galaxy for different final states ($\gamma\gamma$, $W^{+}W^{-}$, $ZZ$, and $hh$), with a solid angle of $1^{\circ}$, compared with the canonical cross-section with a mass of 100 GeV.
+Additionally, we show the total flux predicted at different values of the DM mass.
+We selected the masses to coincide with peaks and local minimums in the SE factor.
+We compare our results with Fermi-LAT sesitivity and the projected CTA sensitivity [21].
+Figure 4: Photons Flux from Annihilation of neutral Dark Matter particles.
+Where it is done for the Navarro-Frenk-White (NFW) profile, with a Line of sight (los) angle in longitude (llos) of $1^{\circ}$, and latitude (blos) of $0^{\circ}$.
+This photon flux is measured at the center of the galaxy, with $\kappa=0$.
+We show differential sensitivity of CTA [21] and Fermi-Lat [22] in the galactic center.
+Figure 5: Photons Flux from Annihilation of neutral Dark Matter particles.
+Where it is done for the Navarro-Frenk-White (NFW) profile, with a Line of sight (los) angle in longitude (llos) of $1^{\circ}$, and latitude (blos) of $0^{\circ}$.
+This photon flux is measured at the center of the galaxy, with $\kappa=1$.
+We show differential sensitivity of CTA [21] and Fermi-Lat [22] in the galactic center.
 5 Summary
-In this paper, we have studied the Sommerfeld enhancement in a model where the DM candidate is a neutral component of a new massive vector field in the fundamnetal representation of $SU(2)_{L}$.
-We found that big resonance in $\sigma v$ for DM masses in the rage of 5 TeV to 10 TeV, in the low velocities regime.
-We used this result to predict the photon flux produced by DM annihilation at the center of the galaxy.
-This prediction can be tested in the near future by CTA.
+In this paper, we have studied the Sommerfeld enhancement in a model where the DM candidate is a neutral component of a new massive vector field in the fundamental representation of $SU(2)_{L}$.
+We found that the annihilation cross-section predicted by our model is generally bigger than the one predicted in the cases of a spin $1/2$ doublet and a spin $1$ triplet.
+Additionaly, the presence of non-minimal interactions of the doublet with the standard gauge model produces a richer resonance structure.
+In fact, in the absence of the non-minimal interaction, only one big resonance in $\sigma v$ is observed for $M_{V}\approx 6$ TeV.
+However, when the non-minimal interactions are present, a second peak appears at $M_{V}\approx 2$ TeV, in the low velocities regime.
+As a product of our work, we used our results on $\sigma v$ to predict the photon flux produced by DM annihilation at the center of the galaxy.
 Acknowledgments
 This work was partially found by Fondecyt grant 1230110, ANID PIA/APOYO AFB220004, and ANID—The Millennium Science Initiative Program ICN2019_044.
-We are very thankful to Motoko Fujiwara for discussions and suggestions that contributed significantly to improve the quality of this paper.
+We are very thankful to Motoko Fujiwara and Patricio Escalona for discussions and suggestions that contributed significantly to improve the quality of this paper.
 Appendix A Electroweak Matrix
@@
 Consequently, the currents in the non-relativistic limit are as follows:
-$J_{A}^{0}(x)J_{A}^{0}(x^{0},\vec{y})=2e^{2}s_{W}^{2}(c_{W}-s_{W}t_{W}^{2})$ $(-[v_{+}^{\dagger}(x)v_{-}^{\dagger}(x^{0},\vec{y})v_{+}(x)v_{-}(x^{0},\vec{y})]$ (118)
-$J_{Z}^{0}(x)J_{Z}^{0}(x^{0},\vec{y})$ $=-g^{2}s_{W}^{2}\theta_{W}[v_{+}^{\dagger}(x)v_{-}^{\dagger}(x^{0},\vec{y})v_{+}(x)v_{-}(x^{0},\vec{y})$
+$J_{A}^{0}(x)J_{A}^{0}(x^{0},\vec{y})=-2g^{2}(c_{W}(1-(\kappa_{1}+\kappa{2})-s_{W}t_{W}^{2})^{2}$ $(-[v_{+}^{\dagger}(x)v_{-}^{\dagger}(x^{0},\vec{y})v_{+}(x)v_{-}(x^{0},\vec{y})]$ (56)
+$J_{Z}^{0}(x)J_{Z}^{0}(x^{0},\vec{y})$ $=-g^{2}(c_{W}(\kappa_{1}+\kappa_{2})+s_{W})^{2}[v_{+}^{\dagger}(x)v_{-}^{\dagger}(x^{0},\vec{y})v_{+}(x)v_{-}(x^{0},\vec{y})$
 $+\left(\dfrac{v_{0}^{\dagger}(x)v_{0}^{\dagger}(x^{0},\vec{y})}{\sqrt{2}}\dfrac{v_{0}(x)v_{0}(x^{0},\vec{y})}{\sqrt{2}}\right)$
@@
 $-\left(v_{0}^{\dagger}(x)a_{0}^{\dagger}(x^{0},\vec{y})v_{0}(x)a_{0}(x^{0},\vec{y})\right)]$ (57)
-$J_{W_{+}}^{0}(x)J_{W_{-}}^{0}(x^{0},\vec{y})=-\dfrac{g}{2}\left[v_{0}^{\dagger}(x)a_{0}^{\dagger}(x^{0},\vec{y})v_{+}(x)v_{-}(x^{0},\vec{y})-v_{-}^{\dagger}(x)v_{+}^{\dagger}(x^{0},\vec{y})a_{0}(x)v_{0}(x^{0},\vec{y})\right]$ (120)
+$J_{W_{+}}^{0}(x)J_{W_{-}}^{0}(x^{0},\vec{y})=$ $-\left[\left(\dfrac{3}{8}-2\sqrt{2}\right)(\kappa_{1}+\kappa_{2})+2\sqrt{2}\right]^{2}$
+$g^{2}\left[v_{0}^{\dagger}(x)a_{0}^{\dagger}(x^{0},\vec{y})v_{+}(x)v_{-}(x^{0},\vec{y})-v_{-}^{\dagger}(x)v_{+}^{\dagger}(x^{0},\vec{y})a_{0}(x)v_{0}(x^{0},\vec{y})\right]$ (58)
 where $s_{W}=\sin\theta_{W}$, $c_{W}=\cos\theta_{W}$ and $t_{W}=\tan\theta_{W}$.
@@
 By plugging these currents in equation (A.1), we obtain:
-$S_{eff,1_{NR}}=-\int d^{4}xd^{3}y~{}s(x,\vec{y})^{\dagger}V_{Gauge}(|\vec{x}-\vec{y}|)s(x,\vec{y})$ (121)
-where $V_{Gauge}(|\vec{x}-\vec{y}|)$ is the electroweak matrix, with $r=|\vec{x}-\vec{y}|$, therefore:
-$V_{Gauge}(r)=-\dfrac{g^{2}}{8\pi r}\begin{pmatrix}4s_{W}^{2}e^{-M_{Z}r}&0&0&0\\ 0&4s_{W}^{2}e^{-M_{Z}r}&0&0\\ 0&0&16s_{W}^{2}e^{-M_{Z}r}&2\sqrt{2}e^{-M_{W}r}\\ 0&0&2\sqrt{2}e^{-M_{W}r}&8s_{W}^{2}e^{-M_{Z}r}-2(c_{W}-s_{W}t_{W}^{2})^{2}\end{pmatrix}$ (126)
+$S_{eff,1_{NR}}=-\int d^{4}xd^{3}y~{}s(x,\vec{y})^{\dagger}V_{g-V}(|\vec{x}-\vec{y}|)s(x,\vec{y})$ (59)
+where $V_{g-V}(|\vec{x}-\vec{y}|)$ is the electroweak matrix, with $r=|\vec{x}-\vec{y}|$, therefore:
+$V_{g-V}(r)=-\dfrac{g^{2}}{8\pi r}\begin{pmatrix}4a^{2}e^{-M_{Z}r}&0&0&0\\ 0&4a^{2}e^{-M_{Z}r}&0&0\\ 0&0&16a^{2}e^{-M_{Z}r}&b^{2}e^{-M_{W}r}\\ 0&0&b^{2}e^{-M_{W}r}&a^{2}e^{-M_{Z}r}-2(c_{W}(1-(\kappa_{1}+\kappa{2})-s_{W}t_{W}^{2})^{2}\end{pmatrix}$ (60)
+with $a=(c_{W}(\kappa_{1}+\kappa_{2})+s_{W})$, $b=(\frac{3}{8}-2\sqrt{2})(\kappa_{1}+\kappa_{2})+2\sqrt{2}$
 A.2 Higgs Boson-Dark Matter
@@
 In order to construct pairs such as $v_{0}v_{0},v_{0}a_{0},a_{0}a_{0}$, and $v_{+}v_{-}$, we need to substitute equations (1) and (36) into (61). Consequently, we obtain:
-$V_{vector}(r)=\dfrac{v^{2}}{32\pi m^{2}r}\begin{pmatrix}8\lambda_{4}^{2}a&0&0&0\\ 0&8\lambda_{4}^{2}a&0&0\\ 0&0&\dfrac{1}{2}(\lambda_{2}+\lambda_{3})^{2}a&(\lambda_{3}+2\lambda_{4})^{2}e^{-M_{W}r}\\ 0&0&(\lambda_{3}+2\lambda_{4})^{2}e^{-M_{W}r}&\lambda_{2}^{2}a\end{pmatrix}$ (132)
-this is electroweak matrix (Higgs Boson-Dark Matter interaction)
+$V_{h-V}(r)=\dfrac{v^{2}}{32\pi M_{V}^{2}r}\begin{pmatrix}8\lambda_{4}^{2}a&0&0&0\\ 0&8\lambda_{4}^{2}a&0&0\\ 0&0&\dfrac{1}{2}(\lambda_{2}+\lambda_{3})^{2}a&(\lambda_{3}+2\lambda_{4})^{2}e^{-M_{W}r}\\ 0&0&(\lambda_{3}+2\lambda_{4})^{2}e^{-M_{W}r}&\lambda_{2}^{2}a\end{pmatrix}$ (62)
+with $a=(2e^{-M_{h}r}+e^{-M_{Z}r})$, this is electroweak matrix (Higgs Boson-Dark Matter interaction)
 Appendix B Annihilation Matrix
@@
 In the framework of effective quantum field theory, these annihilations are depicted by the inclusion of an imaginary potential or an absorptive term in the action, which accounts for the disappearance of pairs of non-relativistic particles.
-These absorptive terms originate from $S{eff,2}$ as outlined in equation (56).
+These absorptive terms originate from $S_{eff,2}$ as outlined in equation (38).
 To provide an interpretation, we can visualize $S_{eff,2}$ in terms of Feynman diagrams.
@@
 $\dfrac{1}{q^{2}+i\epsilon}\rightarrow-2\pi i\delta(q^{2})$ (64)
-Therefore, we perform an integration over the timelike component of the 4-momentum, resulting in the transformation $d^{4}q\delta(q^{2})\rightarrow\frac{d^{3}q}{2|\vec{q}|}$ when the gauge particles are on-shell.
+Therefore, we perform an integration over the time-like component of the 4-momentum, resulting in the transformation $d^{4}q\delta(q^{2})\rightarrow\frac{d^{3}q}{2|\vec{q}|}$ when the gauge particles are on-shell.
 The integration of 4-momentum is then expressed as:
@@
 $\epsilon_{i}(p_{-})\epsilon_{j}(p_{+})\longrightarrow S_{ij}^{J,J_{z}}$ (73)
-with add angular momentun matrices
+with add angular momentum matrices
 $S^{0,0}=-\dfrac{1}{\sqrt{3}}\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix}$ (74)
@@
 $\epsilon_{i}^{\pm}(k_{1})=\dfrac{1}{\sqrt{2}}\begin{pmatrix}\mp\cos\theta\\ -i\\ \pm\sin\theta\end{pmatrix}\hskip 28.45274pt\epsilon_{i}^{\pm}(k_{2})=\dfrac{1}{\sqrt{2}}\begin{pmatrix}\pm\cos\theta\\ -i\\ \mp\sin\theta\end{pmatrix}$ (77)
-The annilihation matrices (Gauge-Dark Matter) are:
-$\Gamma_{WW}^{J=0}=\dfrac{2g^{4}}{64\sqrt{3}\pi m^{2}}\begin{pmatrix}2&0&0&0\\ 0&2&0&0\\ 0&0&1&1\\ 0&0&1&1\end{pmatrix}$ (180)
-$\Gamma_{WW}^{J=2}=\dfrac{g^{4}}{64\pi m^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2&0&0&0\\ 0&2&0&0\\ 0&0&1&1\\ 0&0&1&1\end{pmatrix}$ (185)
-$\Gamma_{AA}^{J=0}=\dfrac{e^{4}}{128\sqrt{3}\pi m^{2}}\begin{pmatrix}2a^{2}&0&0&0\\ 0&2a^{2}&0&0\\ 0&0&a^{2}&4a^{2}\\ 0&0&4a^{2}&4\end{pmatrix}$ (190)
-$\Gamma_{AA}^{J=2}=\dfrac{e^{4}}{256\pi m^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2a^{2}&0&0&0\\ 0&2a^{2}&0&0\\ 0&0&a^{2}&4a^{2}\\ 0&0&4a^{2}&4\end{pmatrix}$ (195)
+The annihilation matrices (Gauge-Dark Matter) are:
+$\Gamma_{WW}^{J=0}=\dfrac{2g^{4}}{64\sqrt{3}\pi M_{V}^{2}}\begin{pmatrix}2&0&0&0\\ 0&2&0&0\\ 0&0&1&1\\ 0&0&1&1\end{pmatrix}$ (78)
+$\Gamma_{WW}^{J=2}=\dfrac{g^{4}}{64\pi M_{V}^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2&0&0&0\\ 0&2&0&0\\ 0&0&1&1\\ 0&0&1&1\end{pmatrix}$ (79)
+$\Gamma_{AA}^{J=0}=\dfrac{e^{4}}{128\sqrt{3}\pi M_{V}^{2}}\begin{pmatrix}2a^{2}&0&0&0\\ 0&2a^{2}&0&0\\ 0&0&a^{2}&4a^{2}\\ 0&0&4a^{2}&4\end{pmatrix}$ (80)
+$\Gamma_{AA}^{J=2}=\dfrac{e^{4}}{256\pi M_{V}^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2a^{2}&0&0&0\\ 0&2a^{2}&0&0\\ 0&0&a^{2}&4a^{2}\\ 0&0&4a^{2}&4\end{pmatrix}$ (81)
 with $a=(1-\tan\theta_{W})$
-$\Gamma_{ZZ}^{J=0}$ $=\dfrac{g^{4}}{128\sqrt{3}\pi m^{2}}\begin{pmatrix}2b^{2}&0&0&0\\ 0&2b^{2}&0&0\\ 0&0&2b^{2}&-4c\\ 0&0&-4c&4\sin^{4}\theta_{W}\end{pmatrix}$ (200)
-$\Gamma_{ZZ}^{J=2}$ $=\dfrac{g^{4}}{256\pi m^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2b^{2}&0&0&0\\ 0&2b^{2}&0&0\\ 0&0&2b^{2}&-4c\\ 0&0&-4c&4\sin^{4}\theta_{W}\end{pmatrix}$ (205)
+$\Gamma_{ZZ}^{J=0}$ $=\dfrac{g^{4}}{128\sqrt{3}\pi M_{V}^{2}}\begin{pmatrix}2b^{2}&0&0&0\\ 0&2b^{2}&0&0\\ 0&0&2b^{2}&-4c\\ 0&0&-4c&4\sin^{4}\theta_{W}\end{pmatrix}$ (82)
+$\Gamma_{ZZ}^{J=2}$ $=\dfrac{g^{4}}{256\pi M_{V}^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2b^{2}&0&0&0\\ 0&2b^{2}&0&0\\ 0&0&2b^{2}&-4c\\ 0&0&-4c&4\sin^{4}\theta_{W}\end{pmatrix}$
 with $b=(\tan^{2}\theta_{W}\sin^{2}\theta_{W}-\sin\theta_{W}\cos\theta_{W})$ y $c=\sin^{2}\theta_{W}\tan\theta_{W}(\sin^{2}\theta_{W}-\cos^{2}\theta_{W})$
-$\Gamma_{AZ}^{J=0}$ $=\dfrac{e^{2}g^{2}}{128\sqrt{3}\pi m^{2}}\begin{pmatrix}d^{2}&0&0&0\\ 0&d^{2}&0&0\\ 0&0&d^{2}&f\\ 0&0&f&\dfrac{\sin^{2}(2\theta_{W})}{\cos^{2}\theta_{W}}\end{pmatrix}$ (210)
-$\Gamma_{AZ}^{J=2}$ $=\dfrac{e^{2}g^{2}}{256\pi m^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}d^{2}&0&0&0\\ 0&d^{2}&0&0\\ 0&0&d^{2}&f\\ 0&0&f&\dfrac{\sin^{2}(2\theta_{W})}{\cos^{2}\theta_{W}}\end{pmatrix}$ (215)
+$\Gamma_{AZ}^{J=0}$ $=\dfrac{e^{2}g^{2}}{128\sqrt{3}\pi M_{V}^{2}}\begin{pmatrix}d^{2}&0&0&0\\ 0&d^{2}&0&0\\ 0&0&d^{2}&f\\ 0&0&f&\dfrac{\sin^{2}(2\theta_{W})}{\cos^{2}\theta_{W}}\end{pmatrix}$ (83)
+$\Gamma_{AZ}^{J=2}$ $=\dfrac{e^{2}g^{2}}{256\pi M_{V}^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}d^{2}&0&0&0\\ 0&d^{2}&0&0\\ 0&0&d^{2}&f\\ 0&0&f&\dfrac{\sin^{2}(2\theta_{W})}{\cos^{2}\theta_{W}}\end{pmatrix}$
 with $d=(\cos\theta_{W}-\sin\theta_{W})$ y $f=(\cos\theta_{W}-\sin\theta_{W})^{4}(1+\tan\theta_{W})^{2}$
@@
 The permissible terms are $V_{0}V_{0},V_{0}V_{0}^{\ast},V_{0}^{\ast}V_{0}^{\ast}$, and $V_{+}V_{-}$, as we require neutral terms. The resulting annihilation matrix is as follows:
-$\Gamma_{vector}=\dfrac{1}{32\pi m^{2}}\begin{pmatrix}\left(\dfrac{3}{2}+\sqrt{2}\right)^{2}\lambda_{4}^{2}&0&0&0\\ 0&\dfrac{9}{4}\lambda_{4}^{2}&0&0\\ 0&0&\dfrac{1}{4}(7\lambda_{2}^{2}+6\lambda_{2}\lambda_{3}+5\lambda_{3}^{2})&\dfrac{1}{2}(\lambda_{3}+2\lambda_{4})^{2}\\ 0&0&\dfrac{1}{2}(\lambda_{3}+2\lambda_{4})^{2}&\dfrac{3}{4}\lambda_{2}^{2}+(\lambda_{2}+\lambda_{3})^{2}\end{pmatrix}$ (221)
+$\Gamma_{h-V}=\dfrac{1}{32\pi M_{V}^{2}}\begin{pmatrix}\left(\dfrac{3}{2}+\sqrt{2}\right)^{2}\lambda_{4}^{2}&0&0&0\\ 0&\dfrac{9}{4}\lambda_{4}^{2}&0&0\\ 0&0&\dfrac{1}{4}(7\lambda_{2}^{2}+6\lambda_{2}\lambda_{3}+5\lambda_{3}^{2})&\dfrac{1}{2}(\lambda_{3}+2\lambda_{4})^{2}\\ 0&0&\dfrac{1}{2}(\lambda_{3}+2\lambda_{4})^{2}&\dfrac{3}{4}\lambda_{2}^{2}+(\lambda_{2}+\lambda_{3})^{2}\end{pmatrix}$ (85)

### v3

@@
 In this way, $\lambda$ will be the only constant regulating the Higgs portal.
+After the electroweak symmetry breaking, the physical mass of $V_{\mu}$ ($M_{V}$) receives contributions from the Proca term ($+m_{V}^{2}V_{\nu}^{\dagger}V^{\nu}$) but also from the interactions with the Higgs field.
 Finally, the $V_{\mu}$ quartic self-interactions will not contribute significantly to the SE since this phenomenon is dominated by interactions mediated by particles much lighter than $V_{\mu}$, like the electroweak gauge bosons and the Higgs boson.
@@
 This effect may produce sensible consequences for the annihilation cross section in the non-relativistic regime [5, 19].
-In the next few paragraphs we will review the formalism used to compute the Sommerfeld enhancement effect and the we will apply it to our model.
-We start by considering the DM annihilation processes, which are described by a two-body state, in the non-relativistic regime.
-The non-relativistic Schwinger-Dyson equation for the Green function $\bra{0}T\phi_{b}(R^{\prime},\textbf{r}^{\prime})\phi_{a}^{\dagger}(R,\textbf{r})\ket{0}$ can be written as:
-$\left[\left(i\partial_{R^{0}}+\frac{\nabla_{R}^{2}}{4M_{V}}+\frac{\nabla_{r}^{2}}{M_{V}}\right)-V(r)+i\frac{N}{2}\Gamma\delta^{3}(r)\right]$ $\bra{0}T\phi_{b}(R^{\prime},\textbf{r}^{\prime})\phi_{a}^{\dagger}(R,\textbf{r})\ket{0}$ (3)
-$=i\delta^{4}(R-R^{\prime})\delta^{3}(\textbf{r}-\textbf{r}^{\prime})\delta_{ab}$
-Here, $m$ represents the mass of Dark Matter, $\phi_{i}$ represents the DM field, $N$ is the number spin states for the two bodies system (for spin 1, we have $N=9$), $V(r)$ is electroweak matrices and $\Gamma$ is the so called annihilation matrix.
-The coordinate $=(R^{0},\textbf{R})$ R ) denotes the center-of-mass coordinate, and r represents the relative coordinate of the Dark Matter multiplets, defined as:
-[…]
-The imaginary part of the diagonal component of this Green function $\bra{0}T\phi_{b}\phi_{a}^{\dagger}\ket{0}$, is associated with the total annihilation cross section.
-We proceed to decompose the Green functions into individual partial wave modes.
-$\bra{0}T\phi_{b}(R^{\prime},\textbf{r}^{\prime})\phi_{a}^{\dagger}(R,\textbf{r})\ket{0}=\int\dfrac{d^{4}P}{(2\pi)^{4}}e^{-iP\cdot(R-R^{\prime})}\sum_{l}\dfrac{2l+1}{4\pi}P_{l}(\cos\gamma)(-i)G_{ba}^{(E,l)}(r,r^{\prime})$ (5)
-where $\gamma$ denotes the relative angle between r and $\textbf{r}^{\prime}$. The non-relativistic energy for a dark matter pair ($E$) is
-$E=P^{0}-\dfrac{\textbf{P}^{2}}{4M_{V}}=\dfrac{M_{V}v^{2}}{4}$ (6)
-where, $v$ represents the relative velocity between a pair of dark matter particles. By substituting (3) into (5), we obtain the following result:
-$\left[-\dfrac{1}{M_{V}r}\dfrac{d^{2}}{dr^{2}}r-\dfrac{l(l+1)}{M_{V}r^{2}}-E+V(r)-i\Gamma\dfrac{N\delta(r)}{8\pi r^{2}}\right]G_{ba}^{(E,l)}(r,r^{\prime})=\dfrac{\delta(r-r^{\prime})}{r^{2}}\delta_{ba}$ (7)
-In this analysis, we focus on the annihilation processes with orbital angular momentum $l=0$ (s-wave), which provide the leading-order contributions in the non-relativistic limit.
-Indeed, we verified that the $l=1$ contribution is at least one order of magnitude smaller than the leading s-wave.
-Consequently, in our calculations the total angular momentum $J$ is equal to the total spin angular momentum $S$.
-The leading-order potential terms exhibit spherical symmetry, resulting in the conservation of $J$ in annihilation processes.
-Furthermore, the conservation of total electric charge $Q$ is upheld.
-As a result, the Green function is expressed in the diagonal form for each eigenstate corresponding to $Q$ and $J$.
-To simplify the differential equation, we introduce $G_{ba}^{(E,l)}(r,r^{\prime})=\frac{g_{ba}(r,r^{\prime})}{rr^{\prime}}$.
-$\left[-\dfrac{1}{M_{V}r}\dfrac{d^{2}}{dr^{2}}r-E+V(r)-i\Gamma\dfrac{N\delta(r)}{8\pi r^{2}}\right]g_{ba}^{(E,l=0)}(r,r^{\prime})=\delta(r-r^{\prime})\delta_{ba}$ (8)
-The cross-section is determined by the imaginary part of the solution, and we derive the leading-order solution within the perturbative expansion of $\alpha_{2}$.
-The leading order solution can be written as
-$g_{ba}(r,r^{\prime})|_{leadingorder}=\dfrac{iNm^{2}}{8\pi}\left[g_{>}(r^{\prime})\Gamma g_{>}^{T}(r)\right]_{ba}$ (9)
-In the above expression, we introduce $g_{>ab}$ which is the solution of the following equation
-$\left[-\dfrac{1}{M_{V}}\dfrac{d}{dr^{2}}-E+V(r)\right]g_{>ab}(r)=0$ (10)
-Where $V(r)=V_{g-V}(r)+V_{h-V}(r)$ represents the electroweak potential, composed of $V_{g-V}(r)$ (interaction gauge bosons-VDM) and $V_{h-V}(r)$ (interaction Higgs boson-VDM) terms, $E$ denotes the energy of dark matter particles, $m$ stands for the mass of dark matter, and $g_{>ab}(r)$ denotes the wave function for $(r\gg 1)$, with the corresponding boundary conditions.
-$g_{>ab}(r=0)=\delta_{ba}$ (11)
-We impose different profiles depending on each potential.
-The short-range Yukawa potential diminishes as $r$ approaches infinity, causing the solution to exhibit an outgoing plane wave behavior.
-In the case of the long-range Coulomb potential, the asymptotic profile is determined by solving the Schrödinger equation in (10).
-In particular, we establish the following condition to govern the asymptotic behavior.
-$g_{>ba}(r)\longrightarrow\ \left\{\begin{array}[]{lcc}d_{ba}(E_{a})e^{i\sqrt{M_{V}E_{b}}r}\hskip 28.45274ptshort-range~{}potential\\ \\ d_{ba}(E_{a})e^{i\sqrt{M_{V}E_{b}}r}e^{i\frac{M_{V}\alpha}{2\sqrt{M_{V}E_{b}}}\log{2\sqrt{M_{V}E_{b}}r}}\hskip 28.45274ptlarge-range~{}potential\\ \end{array}\right.$ (15)
-Here, $E_{a}$ and $E_{b}$ denote the non-relativistic kinetic energy of $\phi_{a}$ and $\phi_{b}$, respectively.
-Given the constraints of numerical calculations, it is not feasible to let $r$ approach infinity.
-Therefore, we enforce these asymptotic behaviors at a sufficiently large value of $r=r_{f}$, ensuring that the numerical solutions become independent of $r_{f}$.
-The deviation from the asymptotic plane wave profile is quantified by $d_{ba}(E_{a})$, which is commonly referred to as the Sommerfeld Factor.
-We numerically solve equation (10) to obtain the Sommerfeld factor as a function of the relative velocity for the initial state $v$.
-This effect becomes significant in the non-relativistic regime ($v\ll 1$), where the potential energy becomes comparable to the kinetic energy ($E\simeq\frac{M_{V}v^{2}}{4}$ ) in the Schrödinger equations.
-Conversely, the enhancement factors approach unity for ($v\approx 1$).
-Consequently, we can recover the perturbative cross-section, as obtained by the Born-Oppenheimer approximation, by replacing the Sommerfeld factor with $1$.
+In the next few paragraphs we will review the formalism used to compute the Sommerfeld enhancement effect [20] and the we will apply it to our model.
 Figure 1: Sommerfeld enhancement effect
-3.2 Annihilation Cross Section
-The optical theorem enables us to establish a connection between $Im(G_{aa}^{E,l=0}(r,r^{\prime}))$ and the s-wave spin-averaged, velocity-weighted annihilation cross-section $\braket{\sigma_{a}v}$.
-This connection results in the following formula.
-[…]
-In the equation above, there is no summation taken over $a$, and $\sigma_{a}v$ represents the annihilation cross section of $\phi_{a}\sim V^{i}V^{j}$.
-We introduce a prefactor denoted as $c_{a}$, defined as follows:
-[…]
-If we substitute the leading order solution given in (9), we obtain
-$\braket{\sigma_{a}v}=c_{a}\dfrac{M_{V}^{2}}{k^{2}}\lim_{E\longrightarrow\frac{k^{2}}{M_{V}^{2}}}\left(E-\dfrac{k^{2}}{m^{2}}\right)^{2}[\mathcal{A}(k)~{}\Gamma~{}\mathcal{A}^{\dagger}(k)]_{aa}$ (21)
-where we define
-[…]
-The asymptotic behavior $r\longrightarrow\infty$ is important to evaluate $\mathcal{A}_{ba}(k)$. Substituting (15) into (22), we obtain
-[…]
-$=d_{ba}(E)\left(-\dfrac{k}{M_{V}}\right)\left(E-\frac{k^{2}}{M_{V}}\right)^{-1}$ (24)
-We replace into (21), we express the annihilation cross section as below
-$\braket{\sigma_{a}v}\simeq c_{a}\sum_{b,a}d_{ab}(E)\Gamma_{bc}d_{ac}^{\ast}(E)$ (25)
-Specifically, we are interested in the velocity-weighted annihilation cross section for the partial annihilation processes $V^{0}V^{0}\longrightarrow XX^{\prime}$, where $X$ represents any Standard Model particle.
-$\braket{\sigma_{a}v}_{XX^{\prime}}=c\sum_{\alpha,\beta}\sum_{J,J_{z}}(\Gamma_{XX^{\prime}}^{J,J_{z}})_{\alpha\beta}~{}d_{3\alpha}(E)d_{3\beta}^{\ast}(E)$ (26)
-Here, the index $3$ must be fixed; therefore, the Sommerfeld enhancement factor implies that we are considering the process $(V_{0}V_{0}^{\ast}\longrightarrow XX^{\prime})$, with $c$ is set to $2$ for initial states consisting of identical particles.
-We define $E$ as approximately $(E\simeq\frac{M_{V}v_{rel}^{2}}{4})$ representing the non-relativistic kinetic energy of the V-particles, $(\Gamma_{XX^{\prime}}^{J,J_{z}})_{\alpha\beta}$ is the Annihilation matrices, with $X$ is any standard model particle, and $d_{\alpha\beta}(E)$ with $(\alpha,\beta=1,2,3,4)$ as the Sommerfeld enhancement factor.
-3.3 Non-Relativistic Limit
+3.1.1 The Two-Particle State and Schrödinger Equation
+To describe the state of two DM-DM particles, which must be neutral in charge and CP-even, we consider the following combinations:
+|$\ket{g}=\begin{pmatrix}v_{0}v_{0}\\ a_{0}a_{0}\\ v_{0}a_{0}\\ v_{+}v_{-}\end{pmatrix}$ (3)
+The order of the components in the vector of equation (3) is arbitrary, and we can use different orderings for the pairs of these doublets.
+Since this effect is non-relativistic, we must use the Schrödinger equation for two bodies, derived from the Schwinger-Dyson equation [5], which is:
+$\left[-\dfrac{1}{M_{V}}\dfrac{d^{2}}{dr^{2}}-\frac{l(l+1)}{M_{V}r^{2}}-E+V(r)-\dfrac{iN\Gamma\delta(r)}{8\pi r^{2}}\right]g_{ab}(r)=0$ (4)
+where index $a$ represents the initial state of the process and $b$ the final state and $r$ is the relative position.
+The wave function $g_{ab}$ describes the two particles in a doublet of $SU(2)_{L}$.
+As we have seen, the potential includes interactions with gauge and Higgs bosons, corresponding to Coulomb (long-range) and Yukawa (short-range) potentials.
+The kinetic energy of these particles is $E\approx\frac{M_{V}v^{2}}{4}$, where $v$ is the relative velocity between them.
+The $\Gamma$ term is the imaginary potential and corresponds to an absorptive part, arising in the low-energy limit of quartic interactions.
+Finally, $N$ is the number of spin states, given by $N=(2l+1)^{2}$, which in our case is 9.
+We can consider polarized initial states in the general case, but the main contribution in the non-relativistic limit occurs in the s-wave, i.e., when $l=0$, where the total angular momentum is given solely by the spin contribution.
+Therefore, the Schrödinger equation is:
+$\left[-\dfrac{1}{M_{V}}\dfrac{d^{2}}{dr^{2}}-E+V(r)-\dfrac{iN\Gamma\delta(r)}{8\pi r^{2}}\right]g_{ab}(r)=0$ (5)
+To solve this equation (5), we need the boundary conditions given by [5]:
+$g_{ab}(r\rightarrow 0)=\delta_{ab}$ (6)
+$g_{ab}(r\rightarrow\infty)=d_{ab}(E)e^{i\sqrt{M_{V}E}r}e^{i\left(\frac{M_{V}\alpha}{2M_{V}E}\ln(2\sqrt{M_{V}E}r)\right)}$ (7)
+where $\delta_{ab}$ is a Kronecker delta, $\alpha$ is the fine-structure constant, and $d_{ab}$ are the Sommerfeld enhancement factors.
+The asymptotic condition of (7) is valid for large distances encompassing the Coulomb and Yukawa potentials.
+The equation (5) can be solved numerically as it does not have an analytical solution.
+Once the solutions are determined, we use the optical theorem, which relates the imaginary part of the Green’s functions to the thermally averaged total annihilation cross-section [5, 7] of DM particles:
+$\braket{\sigma v}_{XX^{\prime}}=c\sum_{a,b}\sum_{J,J_{z}}(\Gamma_{XX^{\prime}}^{J,J_{z}})_{ab}~{}d_{3a}(E)d_{3b}^{\ast}(E)$ (8)
+where $c$ is a normalization constant (2 for $a=b$, 1 for $a\neq b$).
+The fixed index 3 is related to the definition of the vector $\vec{s}$ (14), to select dark matter annihilation $(V_{0}V_{0}\longrightarrow XX^{\prime})$ with $X$ and $X^{\prime}$ being particles of the Standard Model.
+3.1.2 Non-Relativistic Limit
 To ensure that the dark matter vector remains non-relativistic, we must carry out the following non-relativistic limit of the operators, as defined in [7].
@@
 Acknowledgments
-This work was partially found by Fondecyt grant 1230110, ANID PIA/APOYO AFB220004, and ANID—The Millennium Science Initiative Program ICN2019_044.
+This work was partially found by Fondecyt grant 1230110, grant PIIC 2021-II UTFSM, ANID BECA/DOCTORADO NACIONAL 21210616, ANID PIA/APOYO AFB220004, and ANID—The Millennium Science Initiative Program ICN2019_044.
 We are very thankful to Motoko Fujiwara and Patricio Escalona for discussions and suggestions that contributed significantly to improve the quality of this paper.
@@
 A.1 Gauge-Dark Matter
-We are required to calculate the electroweak matrix using the effective action derived from the Lagrangian (2), resulting in the following expression:
-$S_{eff,1}=-\frac{1}{2}\int d^{4}xd^{4}y~{}J^{\mu\dagger}(y)\Delta_{0~{}\mu\nu}^{-1}(x,y)J^{\nu}(x)$ (49)
-Now, we proceed to compute $S_{eff,1}$ as given in equation (49) in the non-relativistic limit, commencing with:
-$\Delta_{0~{}\mu\nu}^{-1}(x,y)=-\int\dfrac{d^{4}q}{(2\pi)^{4}}e^{iq(x-y)}\begin{pmatrix}\frac{1}{q^{2}+i\epsilon}&0&0&0\\ 0&\frac{1}{q^{2}-M_{Z}^{2}+i\epsilon}&0&0\\ 0&0&\frac{1}{q^{2}-M_{W}^{2}+i\epsilon}&0\\ 0&0&0&\frac{1}{q^{2}-M_{W}^{2}+i\epsilon}\end{pmatrix}g_{\mu\nu}$ (50)
-Next, we substitute the gauge currents and (50) into (49), resulting in:
-$S_{eff,1}=\int\dfrac{d^{4}xd^{4}yd^{4}q}{2(2\pi)^{4}}e^{iq(x-y)}\left(\dfrac{J^{\mu}_{A}(x)J_{A~{}\mu}(y)}{q^{2}+i\epsilon}+\dfrac{J^{\mu}_{Z}(x)J_{Z~{}\mu}(y)}{q^{2}-M_{Z}^{2}+i\epsilon}+\dfrac{2J^{\mu}_{W^{+}}(x)J_{W^{-}~{}\mu}(y)}{q^{2}-M_{W}^{2}+i\epsilon}\right)$ (51)
-Now, we need to determine the non-relativistic limit of (51).
-However, we have the relativistic gauge currents expressed as $J_{V}=(J_{V}^{0},\vec{J_{V}}),$ where $V$ represents any gauge boson.
-With this representation, each term of (51) can be expressed as:
-$\dfrac{J^{\mu}_{V}(x)J_{V~{}\mu}(y)}{q^{2}-M_{V}^{2}+i\epsilon}=-\dfrac{J^{0}_{V}(x)J_{V}^{0}(y)}{\vec{q}^{2}+i\epsilon}+\dfrac{J^{0}_{V}(x)J_{V}^{0}(y)q^{0^{2}}}{(\vec{q}^{2}+i\epsilon)(q^{2}-M_{V}^{2}+i\epsilon)}+\dfrac{\vec{J}^{\mu}_{V}(x)\vec{J}_{V\mu}(y)}{q^{2}-M_{V}^{2}+i\epsilon}$ (52)
-The third term in equation (52) involves magnetic interactions between currents.
-To account for non-relativistic currents, we require that $|\vec{J_{V}}|\ll J_{V}^{0}$, and as a result, this term can be neglected.
-The second term is associated with the retarded effects on the interactions between currents.
-Therefore, we must perform integration by parts on this term, leading to:
-$\int$ $d^{4}xd^{4}yd^{4}q~{}e^{iq(x-y)}\dfrac{J_{V}^{0}(x)J_{V}^{0}(y)q^{0^{2}}}{(\vec{q}^{2}+M_{V}^{2})(q^{2}-M_{V}^{2}+i\epsilon)}$
-=$\int$= ∫ $d^{4}xd^{4}yd^{4}q~{}e^{iq(x-y)}\dfrac{\dot{J}_{V}^{0}(x)\dot{J}_{V}^{0}(y)}{(\vec{q}^{2}+M_{V}^{2})(q^{2}-M_{V}^{2}+i\epsilon)}$ (53)
-Using the condition $\mid\frac{1}{v_{i}}\frac{dv_{i}}{dt}\mid\ll M_{V_{0}},$ we can deduce that $|\dot{J}{V}^{0}|\ll J{V}^{0}$, and the retarded effects render this term negligible in the non-relativistic limit.
-Consequently, we need to integrate the first term.
-$\int\dfrac{d^{4}q}{(2\pi)^{4}}\dfrac{e^{iq(x-y)}}{\vec{q}^{2}+M_{V}^{2}}=\delta(x^{0}-y^{0})\int\dfrac{d^{3}q}{(2\pi)^{3}}\dfrac{e^{iq(x-y)}}{\vec{q}^{2}+M_{V}^{2}}=\delta(x^{0}-y^{0})\dfrac{e^{-M_{V}|\vec{x}-\vec{y}|}}{4\pi|\vec{x}-\vec{y}|}$ (54)
-We substitute equation (54) into equation (51), resulting in the following effective action in the non-relativistic limit:
+In the context of non-relativistic dark matter, we need to determine the potentials that interact with these particles in a low-velocity regime.
+For this, we need to determine the real potential (cubic terms) of the Lagrangian (2) or effective action of $SU(2)_{L}$, which is the electroweak matrix, where the conserved currents interact with the gauge fields of the Standard Model, resulting in the following expression:
+$S_{eff,1}\approx\int d^{4}x~{}\left[J_{\mu}^{W^{+}}(x)W^{+\mu}(x)+J_{\mu}^{W^{-}}(x)W^{-\mu}(x)+J_{\mu}^{Z}(x)Z^{\mu}(x)+J_{\mu}^{A}(x)A^{\mu}(x)\right]$ (31)
+After integrating out the gauge bosons in the path integral, we obtain the effective action dependent on the currents.
+We need to calculate the conserved currents from the cubic terms of the Lagrangian (2); the details of the calculation are in  [4].
 $S_{eff,1}=-\int\dfrac{d^{4}xd^{3}y}{8\pi|\vec{x}-\vec{y}|}($ $J_{A}^{0}(x)J_{A}^{0}(x^{0},\vec{y})+J_{Z}^{0}(x)J_{Z}^{0}(x^{0},\vec{y})e^{-M_{Z}|\vec{x}-\vec{y}|}$
 $+2J_{W^{+}}^{0}(x)J_{W^{-}}^{0}(x^{0},\vec{y})e^{-M_{W}|\vec{x}-\vec{y}|})$ (32)
-The first term in equation (A.1) accounts for the Coulomb potential among non-relativistic particles, whereas the remaining terms describe Yukawa potential associated with the exchange of $W$ and $Z$ bosons.
 It is essential to consider only pairs in the form of vectors, as indicated in (14), while other particle combinations are prohibited.
@@
 B.1 Gauge-Dark Matter Interaction
-The non-relativistic vector dark matter particles have the potential to annihilate into relativistic particles, specifically gauge bosons of $SU(2){L}$.
-In the framework of effective quantum field theory, these annihilations are depicted by the inclusion of an imaginary potential or an absorptive term in the action, which accounts for the disappearance of pairs of non-relativistic particles.
-These absorptive terms originate from $S_{eff,2}$ as outlined in equation (38).
+Non-relativistic vector dark matter particles have the potential to annihilate into relativistic particles, specifically gauge bosons of $SU(2)_{L}$.
+In quantum mechanics (Schrödinger equation), these annihilations are represented by an imaginary potential or an absorptive term (quartic terms) in the action, which explains the disappearance of pairs of non-relativistic particles.
+These absorptive terms originate from $S_{eff,2}$.
 To provide an interpretation, we can visualize $S_{eff,2}$ in terms of Feynman diagrams.
-In this representation, the matrix $\Delta_{0}^{-1}$ characterizes the propagators of gauge fields, and each $\mathcal{Q}$ represents a vertex featuring two gauge fields and two vector dark matter fields, as described by Chowdhury et al.
-[4].
-Figure 6: Feynman diagram of quartic terms
+In this representation, the matrix $\Delta$ characterizes the propagators of the gauge fields, and each $\mathcal{U}$ represents a vertex featuring two gauge fields and two vector dark matter fields, as described by Chowdhury et al.
+[4, 24].
 Then, the effective action of absorptive terms is:
-$S_{eff,2}$ $=-iTr\log\left(\Delta_{0}^{-1}\mathcal{Q}\Delta_{0}^{-1}\mathcal{Q}\right)$
-$=-4i\int d^{4}xd^{4}yd^{4}zd^{4}w\dfrac{d^{4}q}{(2\pi)^{4}}\dfrac{d^{4}k}{(2\pi)^{4}}\dfrac{e^{iq(x-y)}}{q^{2}+i\epsilon}\dfrac{e^{ik(w-z)}}{k^{2}+i\epsilon}Tr\left(\mathcal{Q}(y,z)\mathcal{Q}(w,x)\right)$ (63)
-The factor of 4 corresponds to the trace of gauge fields’ indices, and our focus is on the imaginary part. To proceed, we consider the gauge bosons in the loop to be on-shell, as follows:
-$\dfrac{1}{q^{2}+i\epsilon}\rightarrow-2\pi i\delta(q^{2})$ (64)
-Therefore, we perform an integration over the time-like component of the 4-momentum, resulting in the transformation $d^{4}q\delta(q^{2})\rightarrow\frac{d^{3}q}{2|\vec{q}|}$ when the gauge particles are on-shell.
-The integration of 4-momentum is then expressed as:
-$\int\dfrac{d^{4}q}{(2\pi)^{4}}\dfrac{d^{4}k}{(2\pi)^{4}}\dfrac{e^{iq(x-y)}}{q^{2}+i\epsilon}\dfrac{e^{ik(w-z)}}{k^{2}+i\epsilon}=-\dfrac{i}{8\pi}\delta^{4}(x-y)$ (65)
-Then, replace into equation (B.1), we can obtain:
-$S_{eff,2}=\frac{i}{2\pi}\int d^{4}xTr\begin{pmatrix}Q_{AA}(x)&Q_{AZ}(x)&Q_{AW}(x)&Q_{AW}^{\ast}(x)\\ Q_{AZ}(x)&Q_{ZZ}(x)&Q_{ZW}(x)&Q_{ZW}^{\ast}(x)\\ Q_{AW}^{\ast}(x)&Q_{ZW}^{\ast}(x)&Q_{WW}(x)&0\\ Q_{AW}^{\ast}(x)&Q_{ZW}^{\ast}(x)&0&Q_{WW}^{\ast}(x)\\ \end{pmatrix}^{2}$ (66)
+$S_{eff,2}$ $=-iTr\log\left(\Delta^{-1}\mathcal{U}\Delta^{-1}\mathcal{U}\right)$
+$=\frac{i}{2\pi}\int d^{4}x~{}Tr\left[\begin{pmatrix}U_{AA}(x)&U_{AZ}(x)&U_{AW}(x)&U_{AW}^{\ast}(x)\\ U_{AZ}(x)&U_{ZZ}(x)&U_{ZW}(x)&U_{ZW}^{\ast}(x)\\ U_{AW}^{\ast}(x)&U_{ZW}^{\ast}(x)&U_{WW}(x)&0\\ U_{AW}^{\ast}(x)&U_{ZW}^{\ast}(x)&0&U_{WW}^{\ast}(x)\\ \end{pmatrix}\right]^{2}$ (40)
 We need only pairs like $V_{0}V_{0},V_{0}^{\ast}V_{0}^{\ast},V_{0}V_{0}^{\ast}$ and $V_{+}V_{-}$, then
-$S_{eff,2}$ $=\dfrac{i}{2\pi}\int~{}d^{4}x\left(Q_{AA}(x)^{2}+Q_{ZZ}(x)^{2}+2Q_{WW}(x)^{2}+2Q_{AZ}(x)^{2}\right)$
+$S_{eff,2}$ $=\dfrac{i}{2\pi}\int~{}d^{4}x\left(U_{AA}(x)^{2}+U_{ZZ}(x)^{2}+2U_{WW}(x)^{2}+2U_{AZ}(x)^{2}\right)$
 $=-\int~{}d^{4}x\dfrac{g^{2}}{8}[2W_{\mu}^{+}V_{\nu}^{-}V^{+~{}\nu}W^{-~{}\mu}+2W_{\mu}^{-}(V_{\nu}^{0})^{\ast}V^{0~{}\nu}W^{+~{}\mu}$
@@
 The Q’s terms are:
-$Q_{WW}^{2}(x)$ $=\dfrac{g^{4}}{16}(V_{\nu}^{-}V^{+~{}\nu}+(V_{\nu}^{0})^{\ast}V^{0~{}\nu})^{2}$ (69)
+$U_{WW}^{2}(x)$ $=\dfrac{g^{4}}{16}(V_{\nu}^{-}V^{+~{}\nu}+(V_{\nu}^{0})^{\ast}V^{0~{}\nu})^{2}$ (43)
 $=\dfrac{g^{4}}{16}(V_{i}^{-}V^{+~{}j}V_{i}^{-}V^{+~{}j}+(V_{i}^{0})^{\ast}V^{0~{}j}(V_{i}^{0})^{\ast}V^{0~{}j}$
 $+2(V_{i}^{0})^{\ast}V^{0~{}j}(V_{i}^{-})^{\ast}V^{+~{}j})$
-$Q_{AA}^{2}(x)=\dfrac{e^{4}}{64}\left[(1-\tan\theta_{W})^{2}((V_{\nu}^{0})^{\ast}V^{0~{}\nu})^{2}+4(V_{\nu}^{-}V^{+~{}\nu})^{2}+4(1-\tan\theta_{W})(V_{\nu}^{0})^{\ast}V^{0~{}\nu}V_{\nu}^{-}V^{+~{}\nu}\right]$ (70)
-$Q_{ZZ}^{2}(x)=\dfrac{g^{4}}{64}$ $[(\tan^{2}\theta_{W}\sin^{2}\theta_{W}-\sin\theta_{W}\cos\theta_{W})^{2}((V_{\nu}^{0})^{\ast}V^{0~{}\nu})^{2}+4\sin^{4}\theta_{W}(V_{\nu}^{-}V^{+~{}\nu})^{2}$
+$U_{AA}^{2}(x)=\dfrac{e^{4}}{64}\left[(1-\tan\theta_{W})^{2}((V_{\nu}^{0})^{\ast}V^{0~{}\nu})^{2}+4(V_{\nu}^{-}V^{+~{}\nu})^{2}+4(1-\tan\theta_{W})(V_{\nu}^{0})^{\ast}V^{0~{}\nu}V_{\nu}^{-}V^{+~{}\nu}\right]$ (44)
+$U_{ZZ}^{2}(x)=\dfrac{g^{4}}{64}$ $[(\tan^{2}\theta_{W}\sin^{2}\theta_{W}-\sin\theta_{W}\cos\theta_{W})^{2}((V_{\nu}^{0})^{\ast}V^{0~{}\nu})^{2}+4\sin^{4}\theta_{W}(V_{\nu}^{-}V^{+~{}\nu})^{2}$
 $-4\sin^{2}\theta_{W}(\tan^{2}\theta_{W}\sin^{2}\theta_{W}-\sin\theta_{W}\cos\theta_{W})^{2}(V_{\nu}^{0})^{\ast}V^{0~{}\nu}V_{\nu}^{-}V^{+~{}\nu}]$ (45)
-$Q_{AZ}^{2}(x)$ $=\dfrac{e^{2}g^{2}}{64}[(\cos\theta_{W}-\sin\theta_{W})^{2}((V_{\nu}^{0})^{\ast}V^{0~{}\nu})^{2}+\dfrac{\cos^{2}(2\theta_{W})}{\cos^{2}\theta_{W}}{\cos\theta_{W}}(V_{\nu}^{-}V^{+~{}\nu})^{2}$
+$U_{AZ}^{2}(x)$ $=\dfrac{e^{2}g^{2}}{64}[(\cos\theta_{W}-\sin\theta_{W})^{2}((V_{\nu}^{0})^{\ast}V^{0~{}\nu})^{2}+\dfrac{\cos^{2}(2\theta_{W})}{\cos^{2}\theta_{W}}{\cos\theta_{W}}(V_{\nu}^{-}V^{+~{}\nu})^{2}$
 $+(\cos\theta_{W}-\sin\theta_{W})^{4}(1+\tan\theta_{W})^{2}(V_{\nu}^{0})^{\ast}V^{0~{}\nu}V_{\nu}^{-}V^{+~{}\nu}]$ (46)
@@
 $\epsilon_{i}(p_{-})\epsilon_{j}(p_{+})\longrightarrow S_{ij}^{J,J_{z}}$ (47)
-with add angular momentum matrices
-$S^{0,0}=-\dfrac{1}{\sqrt{3}}\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix}$ (74)
-$S^{1,1}=\dfrac{1}{2}\begin{pmatrix}0&0&-1\\ 0&0&-i\\ 1&i&0\end{pmatrix}\hskip 28.45274ptS^{1,0}=\dfrac{1}{\sqrt{2}}\begin{pmatrix}0&i&0\\ -i&0&0\\ 0&0&0\end{pmatrix}\hskip 28.45274ptS^{1,-1}=\dfrac{1}{2}\begin{pmatrix}0&0&-1\\ 0&0&i\\ 1&-i&0\end{pmatrix}$ (75)
-$S^{2,2}=\dfrac{1}{2}\begin{pmatrix}1&i&0\\ i&-1&0\\ 0&0&0\end{pmatrix}\hskip 28.45274ptS^{2,1}=\dfrac{1}{2}\begin{pmatrix}0&0&-1\\ 0&0&-i\\ -1&-i&0\end{pmatrix}\hskip 28.45274ptS^{2,0}=\dfrac{1}{\sqrt{6}}\begin{pmatrix}-1&0&0\\ 0&-1&0\\ 0&0&2\end{pmatrix}$
-$S^{2,-1}=\dfrac{1}{2}\begin{pmatrix}0&0&1\\ 0&0&-i\\ 1&-i&0\end{pmatrix}\hskip 28.45274ptS^{2,-2}=\dfrac{1}{2}\begin{pmatrix}1&-i&0\\ -i&-1&0\\ 0&0&0\end{pmatrix}$ (76)
-We are using polarization vector for photons as:
-$\epsilon_{i}^{\pm}(k_{1})=\dfrac{1}{\sqrt{2}}\begin{pmatrix}\mp\cos\theta\\ -i\\ \pm\sin\theta\end{pmatrix}\hskip 28.45274pt\epsilon_{i}^{\pm}(k_{2})=\dfrac{1}{\sqrt{2}}\begin{pmatrix}\pm\cos\theta\\ -i\\ \mp\sin\theta\end{pmatrix}$ (77)
 The annihilation matrices (Gauge-Dark Matter) are:

### v4

@@
 Abstract
-In this work, we study an extension of the Standard Model which include a new massive vector field in the fundamental representation of $SU(2)_{L}$.
+In this work, we study an extension of the Standard Model which includes a new massive vector field in the fundamental representation of $SU(2)_{L}$.
 The neutral component of the new vector field is a natural dark matter candidate.
 We compute the Annihilation Cross Section with Sommerfeld Enhancement and the Photon Flux arising from Dark Matter annihilation events.
-We found that the model exhibits a resonant mass part around the range of $2-10$ TeV.
+We found that the model exhibits resonances for masses in the range of $2-10$ TeV.
 This resonance is influenced by the choice of portal Higgs coupling parameter.
@@
 An intriguing realization of the WIMP idea is to postulate the DM as the neutral component of an electroweak multiplet.
-This paradigm, usually called “Minimal Dark Matter”, was proposed long ago [3] and have been studied in detail for scalar field [4] and fermionic fields [5].
-In the case of vector DM (VDM), an effective model including a VDM candidate originated from a vector field in the adjoint representation of $SU(2)_{L}$ was first proposed in [6] while an ultraviolet complete model was developed in [7].
+This paradigm, usually called “Minimal Dark Matter”, was proposed long ago [3] and has been studied in detail for scalar fields [4] and fermionic fields [5].
+In the case of vector DM (VDM), an effective model that included a candidate VDM originated from a vector field in the adjoint representation of $SU(2)_{L}$ was first proposed in [6], while an complete ultraviolet model was developed in [7].
 On the other hand, models including a vector DM candidate coming from an electroweak doublet were studied in  [8, 9].
-These vector models are consistent with experimental data, if the new fields has mass in the range of some few TeV.
-In this mass range, direct detection [10] becomes difficult and the discovery at colliders seems to be challenging even for future colliders [6].
-In this circumstances, indirect search can be a privileged alternative to discover or rule out these models.
+These vector models are consistent with experimental data, if the new fields have a mass in the range of some few TeV.
+In this mass range, direct detection [10] becomes difficult, and discovery in colliders seems to be challenging even for future colliders [6].
+In these circumstances, indirect search can be a preferred alternative to discover or rule out these models.
 However, it is well known that to obtain an accurate value of the DM annihilation cross section in the non-relativistic regime, it is necessary to take into account non-perturbative effects like the so called Sommerfeld Enhancement (SE).
-The SE effect was computed for the cases of scalar DM [4, 11, 12, 13, 14], fermion DM field [5, 15] in the doublet and triplet representation of $SU(2)_{L}$ and vector DM in the adjoint representation  [7].
-In this work, we focus on computing the SE for the annihilation of VDM in the fundamental representation of $SU(2)_{L}$ and the photon flux (relevant for Indirect Detection  [10, 16, 17, 18]) produced by pair annihilation of the VDM.
-As we will see in the following section, a unique feature of our case is the presence of non-minimal interactions between the massive vector doublet and the standard gauge bosons.
+The consequences of SE have been extensively studied in various scenarios, including scalar dark matter (DM)[4, 11, 12, 13, 14], fermionic DM fields[5, 15] in the doublet and triplet representations of $SU(2)_{L}$, and vector DM in the adjoint representation [7].
+In this work, we focus on computing the SE for the annihilation of vector DM (VDM) in the fundamental representation of $SU(2)_{L}$ and analyzing the resulting photon flux, which is relevant for indirect detection [10, 16, 17, 18].
+A distinctive feature of our model is the presence of non-minimal interactions between the massive vector doublet and the standard gauge bosons, which introduce novel dynamics in the annihilation process.
+Moreover, because the DM particles in our scenario are significantly heavier than the $W$ and $Z$ bosons, the SE effect becomes particularly important and impactful.
 Our work is organized as follows.
 In section 2, we describe the main aspects of the model.
-Section 3 is devoted to the computation of the electroweak, annihilation matrices and annihilation cross section with SE.
+Section 3 is devoted to the computation of the electroweak, annihilation matrices, and annihilation cross section with SE.
 In section 4 we apply our results to compute the photon flux from DM annihilation at the center of the Milky Way and describe our results.
@@
 Similar non-minimal interactions in the scalar and fermion cases happen to be higher dimensional.
-In order to avoid a direct photon interaction with $V_{\mu}^{0}$, the otherwise arbitrary constants $\kappa_{1}$ and $\kappa_{2}$ have to be equal.
+In order to avoid a direct photon interaction with $V_{\mu}^{0}$, the otherwise arbitrary constants $\kappa_{1}$ and $\kappa_{2}$ have to be equal (where $\kappa_{1}$ and $\kappa_{2}$ are parameters analogous to couplings in the context of leptoquark vector models).
 For the purpose of comparing, we work with two values of $\kappa=\kappa_{1}=\kappa_{2}$: $\kappa=0$ and $\kappa=1$.
 The last value coincides with the choice made in [8].
+This model is a simple extension that can also explain the mass of neutrinos, which are not included in the SM.
 Another important sector of the Lagrangian is the part of the potential which describes the interaction between the Higgs field and our massive vector doublet.
 Of course, the coupling constant that regulates the interactions between the Higgs field and the new vector field are free parameters.
-However, there is no reason to think that an important hierarchy arise among these couplings.
+However, there is no reason to think that an important hierarchy arises among these couplings.
 Consequently, and for the sake of simplicity, we will take the assumption that all of them are of the same order $\lambda_{2}\approx\lambda_{3}\approx\lambda_{4}\approx\lambda$.
@@
 Consequently, this accidental $Z_{2}$ symmetry makes the lighter component of $V_{\mu}$ stable.
-This stable component happens to be the neutral one ($V_{\mu}^{0}$)
+This stable component happens to be the neutral one ($V_{\mu}^{0}$).
 3 Sommerfeld Enhancement Computation
 3.1 Basic Formalism
-The so called “Sommerfeld enhancement” is a non-perturbative effect due to the resumation of initial states interactions.
-This effect may produce sensible consequences for the annihilation cross section in the non-relativistic regime [5, 19].
-In the next few paragraphs we will review the formalism used to compute the Sommerfeld enhancement effect [20] and the we will apply it to our model.
+We now compute the indirect signals of $\gamma$-rays produced by dark matter (DM) pair annihilation in specific astrophysical environments, with a particular focus on the Galactic Center of the Milky Way.
+Our analysis includes the non-relativistic effect known as the Sommerfeld enhancement, which accounts for the resummation of interactions between initial states.
+This effect can significantly amplify the annihilation cross-section, making it a crucial factor in indirect DM searches [5, 19].
+Recent observational data from Fermi LAT and CTA have provided stringent constraints on $\gamma$-ray signals associated with DM annihilation in the Galactic Center, offering invaluable insights into DM properties and annihilation cross-sections [20].
+These observations underscore the importance of incorporating the Sommerfeld enhancement in theoretical models to accurately predict $\gamma$-ray fluxes.
+In the following sections, we review the formalism employed to calculate the Sommerfeld enhancement effect [21].
+We then apply this framework to our model, evaluating its implications for indirect detection strategies using $\gamma$-ray observatories such as Fermi LAT and CTA.
 Figure 1: Sommerfeld enhancement effect
@@
 $\left[-\dfrac{1}{M_{V}}\dfrac{d^{2}}{dr^{2}}-\frac{l(l+1)}{M_{V}r^{2}}-E+V(r)-\dfrac{iN\Gamma\delta(r)}{8\pi r^{2}}\right]g_{ab}(r)=0$ (4)
-where index $a$ represents the initial state of the process and $b$ the final state and $r$ is the relative position.
-The wave function $g_{ab}$ describes the two particles in a doublet of $SU(2)_{L}$.
-As we have seen, the potential includes interactions with gauge and Higgs bosons, corresponding to Coulomb (long-range) and Yukawa (short-range) potentials.
-The kinetic energy of these particles is $E\approx\frac{M_{V}v^{2}}{4}$, where $v$ is the relative velocity between them.
-The $\Gamma$ term is the imaginary potential and corresponds to an absorptive part, arising in the low-energy limit of quartic interactions.
-Finally, $N$ is the number of spin states, given by $N=(2l+1)^{2}$, which in our case is 9.
+The index $a$ represents the initial state of the process, while $b$ corresponds to the final state, and $r$ denotes the relative position.
+The wave function $g_{ab}(r)$ describes the two particles within a doublet of $SU(2)_{L}$.
+As previously discussed, the potential $V(r)$ originates from trilinear interactions between the new vector bosons and the electroweak bosons, including the Higgs boson.
+This can be expressed as $V(r)=V_{gauge}(r)+V_{Higgs}(r)$, incorporating Coulomb (long-range) and Yukawa (short-range) interactions.
+The kinetic energy of these particles is approximately $E\approx\frac{M_{V}v^{2}}{2}$, where vv is their relative velocity.
+The term $\Gamma$ represents the imaginary part of the potential and corresponds to an absorptive component arising from quartic interactions in the low-energy limit.
+Lastly, $N$ denotes the number of spin states, given by $N=(2l+1)$, which in our specific case evaluates to 9.
 We can consider polarized initial states in the general case, but the main contribution in the non-relativistic limit occurs in the s-wave, i.e., when $l=0$, where the total angular momentum is given solely by the spin contribution.
@@
 $\braket{\sigma v}=\braket{\sigma v}_{W^{+}W^{-}}+\braket{\sigma v}_{ZZ}+\braket{\sigma v}_{\gamma\gamma}+\braket{\sigma v}_{Z\gamma}+\braket{\sigma v}_{hh}.$ (28)
-Our results are shown in figures 2 and 3 .
-The first one shows our predictions without taking into account the non-minimal interactions ($\kappa=0$) while in the second figure we used $\kappa=1$.
-In both cases, there are several noteworthy characteristics of the Sommerfeld enhancement that merit discussion.
-Firstly, as usual, the SE is a velocity-dependent phenomenon.
-For instance, no structure is observed for a relative velocity of $v=0.3$ (dashed lines in figures 2 and 3), which is representative of the early Universe, when the relic density was formed.
-In contrast, for smaller velocities, like the one expected in the galaxy ($v=10^{-3}$), significant oscillations and resonances appear.
-We have observed that the small oscillations dependent on interactions mediated by the Higgs boson.
-Indeed, we see that they are attenuated for large values of $\lambda$ parameter Interestingly, the presence of the non-minimal interactions introduce an additional resonance at a lower mass.
-However, the $\kappa$-terms don’t modify significantly the size of the oscillations nor the amplitude of the resonances.
-For masses that exhibit such resonances (roughly around $1.9$ and $6$ TeV), the Sommerfeld factor experiences a significant boost of about 5 order of magnitude, for all the values of portal coupling constant $\lambda$.
+Our results are presented in Figures 2 and 3.
+The first figure displays our predictions without considering the non-minimal interactions $(\kappa=0)$, while the second figure incorporates these interactions with $(\kappa=1)$.
+In both cases, several notable features of the Sommerfeld enhancement (SE) emerge that warrant discussion.
+Firstly, as is well-known, SE is a velocity-dependent phenomenon.
+For instance, at a relative velocity of $(v=0.3)$ (dashed lines in Figures 2 and 3), which corresponds to the conditions of the early Universe when the relic density was established, no significant structure is observed.
+This is because the particles in this regime possess relativistically significant velocities, and the SE, being a non-relativistic effect, is not impacted under such conditions.
+In contrast, for smaller velocities such as those expected within the galaxy $(v=10^{-3})$, pronounced oscillations and resonances become evident.
+We observed that the small oscillations depend on interactions mediated by the Higgs boson.
+Notably, these oscillations are attenuated for large values of the portal coupling constant $\lambda$.
+Interestingly, the inclusion of non-minimal interactions $(\kappa\neq 0$ introduces an additional resonance at a lower mass.
+However, the $\kappa$-terms do not significantly alter the magnitude of the oscillations or the amplitude of the resonances.
+For masses that exhibit such resonances, approximately around $1.91$ TeV and $6$ TeV, the Sommerfeld factor experiences a substantial boost, increasing by nearly five orders of magnitude across all values of the portal coupling constant $\lambda$.
+This remarkable enhancement underscores the critical role of these parameters in shaping the indirect detection prospects for dark matter.
 Figure 2: Annihilation cross-section (fundamental representation spin-1 of $SU(2)_{L}$) for the process $v_{0}a_{0}\longrightarrow SMSM$ with the Sommerfeld Enhancement factor, considering velocities $v=10^{-3}$ (galactic Halo, shown as continuous lines, green, blue, black and gray lines ) and $v=0.3$ (Relic density, indicated by dashed lines), and different values of the Higgs portal coupling constant $\lambda$, with $\kappa_{1}=\kappa_{2}=0$.
@@
 Here, $\braket{\sigma v}_{XX^{\prime}}$ corresponds to the total cross-section for Dark Matter annihilation, as shown in equation (8).
-Using the expressions above, we calculate the photon flux resulting from the annihilation cross section (8), which serves as a method for the indirect detection of DM.
+Using the expressions above, we calculated the photon flux resulting from the annihilation cross section (8) using the manual from [23], which serves as a method for the indirect detection of DM.
 This entails the integration of equation (29).
@@
 In all the cases we have used $\lambda=0.5$.
-We present the signals of photon flux resulting from the annihilation of VDM particles from the center of the galaxy for different final states ($\gamma\gamma$, $W^{+}W^{-}$, $ZZ$, and $hh$), with a solid angle of $1^{\circ}$, compared with the canonical cross-section with a mass of 100 GeV.
-Additionally, we show the total flux predicted at different values of the DM mass.
+We present the signals of photon flux resulting from the annihilation of VDM particles from the center of the galaxy for different final states ($\gamma\gamma$, $W^{+}W^{-}$, $ZZ$, and $hh$), with a solid angle of $1^{\circ}$, compared with the canonical cross-section with a mass of 100 GeV.Additionally, we show the total flux predicted at different values of the DM mass.
 We selected the masses to coincide with peaks and local minimums in the SE factor.
@@
 Figure 4: Photons Flux from Annihilation of neutral Dark Matter particles.
-Where it is done for the Navarro-Frenk-White (NFW) profile, with a Line of sight (los) angle in longitude (llos) of $1^{\circ}$, and latitude (blos) of $0^{\circ}$.
+Where it is done for the Navarro-Frenk-White (NFW) profile, considering a region of the center of the Milky Way with solid angle $\Delta\Omega=0.96\times 10^{-3}$ sr (see Table 2 in [24]).
+This photon flux is measured at the center of the galaxy.
 This photon flux is measured at the center of the galaxy, with $\kappa=0$.
@@
 Figure 5: Photons Flux from Annihilation of neutral Dark Matter particles.
-Where it is done for the Navarro-Frenk-White (NFW) profile, with a Line of sight (los) angle in longitude (llos) of $1^{\circ}$, and latitude (blos) of $0^{\circ}$.
+Where it is done for the Navarro-Frenk-White (NFW) profile, considering a region of the center of the Milky Way with solid angle $\Delta\Omega=0.96\times 10^{-3}$ sr (see Table 2 in [24]).
 This photon flux is measured at the center of the galaxy, with $\kappa=1$.
@@
 We found that the annihilation cross-section predicted by our model is generally bigger than the one predicted in the cases of a spin $1/2$ doublet and a spin $1$ triplet.
-Additionaly, the presence of non-minimal interactions of the doublet with the standard gauge model produces a richer resonance structure.
-In fact, in the absence of the non-minimal interaction, only one big resonance in $\sigma v$ is observed for $M_{V}\approx 6$ TeV.
-However, when the non-minimal interactions are present, a second peak appears at $M_{V}\approx 2$ TeV, in the low velocities regime.
-As a product of our work, we used our results on $\sigma v$ to predict the photon flux produced by DM annihilation at the center of the galaxy.
+Additionally, the presence of non-minimal interactions of the doublet with the standard gauge model produces a richer resonance structure.
+Our results show that certain regions of the parameter space of our mode are testable by facilities like CTA.
 Acknowledgments